### v2

# FuxiTranyu: A Multilingual Large Language Model Trained with Balanced Data

Abstract
Large language models (LLMs) have demonstrated prowess in a wide range of tasks. However, many LLMs exhibit significant performance discrepancies between high- and low-resource languages. To mitigate this challenge, we present FuxiTranyu, an open-source multilingual LLM, which is designed to satisfy the need of the research community for balanced and high-performing multilingual capabilities. FuxiTranyu-8B, the base model with 8 billion parameters, is trained from scratch on a meticulously balanced multilingual data repository that contains 600 billion tokens covering 43 natural languages and 16 programming languages. In addition to the base model, we also develop two instruction-tuned models: FuxiTranyu-8B-SFT that is fine-tuned on a diverse multilingual instruction dataset, and FuxiTranyu-8B-DPO that is further refined with DPO on a preference dataset for enhanced alignment ability. Extensive experiments on a wide range of multilingual benchmarks demonstrate the competitive performance of FuxiTranyu against existing multilingual LLMs, e.g., BLOOM-7B, PolyLM-13B, Llama-2-Chat-7B and Mistral-7B-Instruct. Interpretability analyses at both the neuron and representation level suggest that FuxiTranyu is able to learn consistent multilingual representations across different languages. To promote further research into multilingual LLMs and their working mechanisms, we release both the base and instruction-tuned FuxiTranyu models together with 58 pretraining checkpoints at HuggingFace [^1] https://huggingface.co/TJUNLP/FuxiTranyu-8B and Github. [^2] https://github.com/tjunlp-lab/FuxiTranyu
11footnotetext: Correspondence to: Deyi Xiong.
1 Introduction
A well-pretrained base model plays a pivotal role in facilitating research and applications of large language models. However, training a base LLM from scratch typically demands a substantial amount of data and significant computational resources, posing a barrier to the development of new LLMs. On the other hand, the majority of LLMs are usually tailored to specific languages such as English Touvron et al. (2023a, b) or Chinese Bai et al. (2023), neglecting the high demand for multilingual capabilities across multiple languages, especially low-resource languages. While certain LLMs, such as Mistral models Jiang et al. (2023a), demonstrate multilingual capabilities, their language coverage remains limited. This limitation significantly restricts the exploration of multilingualism in LLMs under the massive multilingual setting.
LLMs Pre-training Tokens Languages
Base
Model Available
Pretraining
Checkpoints Available
BLOOM-7B1 Scao et al. (2022a) 300B 46 NLs + 13 PLs $\times$
Aya 23-8B Aryabumi et al. (2024) Unknown 23 NLs $\times$ $\times$
PolyLM-13B Wei et al. (2023) 638B 18 NLs $\times$
FuxiTranyu-8B 606B 43 NLs + 16 PLs
Table 1: Comparison between trending multilingual large language models and FuxiTranyu, where NL stands for natural language while PL for programming language.
Recent efforts have been dedicated towards mitigating such language-specific constraints through supervised fine-tuning, as exemplified by Okapi Lai et al. (2023). However, as highlighted by the alignment hypothesis in LIMA Zhou et al. (2024), the knowledge and capabilities of LLMs are predominantly derived from pre-training rather than supervised fine-tuning. Supervised fine-tuning primarily serves to align the behaviors of these models with instructions, which constitutes a sub-distribution of the pre-training data. Consequently, for LLMs whose pre-training data are dominated by a few languages, the effectiveness of supervised fine-tuning in enhancing their multilingual capabilities might be limited.
Other initiatives have focused on pre-training multilingual LLMs, such as BLOOM Scao et al. (2022a) and PolyLM Wei et al. (2023). Nevertheless, these efforts are hindered by their performance, which does not measure up to that of current trending LLMs. BLOOM suffers from outdated training data while PolyLM is undermined by imbalanced language distribution, with English data accounting for approximately 70% and Chinese for ~20%, potentially leading to insufficient learning of under-represented languages. Previous studies Xu et al. (2024) disclose three traits of multilingual LLMs caused by imbalanced language resources in pre-training: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, suggesting that multilingual LLMs could benefit from balanced data distribution across languages.
Recently introduced multilingual LLMs, e.g., Aya 23 models Aryabumi et al. (2024), have demonstrated remarkable performance on multiple multilingual benchmarks. They are derived from the CommandR series of models [^3] https://cohere.com/command by performing supervised fine-tuning. However, only the weights of Aya 23 have been released, with its base model remaining undisclosed.
In this work, we present FuxiTranyu, a family of multilingual LLMs supporting 43 natural languages and 16 programming languages. The FuxiTranyu initiative aims to mitigate the aforementioned challenges of multilingual LLMs. The base model comprises 8 billion parameters and has been trained from scratch using approximately 600 billion multilingual tokens. To ensure balanced learning across all supported languages, we have manually controlled the sampling ratio of pre-training data for different languages, striving for as balanced distribution as possible. In line with our commitment to advancing research in multilingual LLMs, we have also released 58 pre-training checkpoints, resonating with the efforts of LLM360 Liu et al. (2023). Table 1 compares FuxiTranyu with currently available multilingual LLMs from different perspectives.
In addition to the base model, we develop two instruction-tuned models, FuxiTranyu-8B-SFT that is fine-tuned on a collected high-quality multilingual instruction dataset, and FuxiTranyu-8B-DPO that is further tuned on preferences with DPO for enhanced alignment ability.
To evaluate multilingual capabilities of the FuxiTranyu models, we have conducted extensive evaluations across multiple domains, encompassing multilingual discriminative tasks such as multilingual ARC, HellaSwag, and MMLU Lai et al. (2023), XWinograd Muennighoff et al. (2022); Tikhonov and Ryabinin (2021), XCOPA Ponti et al. (2020), XStoryCloze Lin et al. (2021), and multilingual generative tasks including WMT and IWSLT translation benchmarks Bojar et al. (2016); Cettolo et al. (2017) and XL-Sum summarization benchmark Hasan et al. (2021). Our evaluations focus on knowledge, capability and alignment dimensions categorized by Guo et al. (2023). As detailed in Section 5, FuxiTranyu models have demonstrated superior performance on the multilingual ARC, HellaSwag, MMLU, XWinograd, XCOPA, and XStoryCloze compared to BLOOM-7B1 and PolyLM-13B. Furthermore, our two instruction-tuned models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, outperform Llama-2-Chat-7B, Mistral-7B-Instruct-v0.1, BLOOMZ-7B1, PolyLM-MultiAlpaca-13B on translation benchmarks. FuxiTranyu also achieves remarkable results on summarization.
To provide a deep understanding of the multilingual capabilities of FuxiTranyu models, we have conducted interpretability analyses from two distinct perspectives: neuron analysis and representation analysis, as detailed in Section 6. Analysis results indicate that FuxiTranyu-8B has learned more language-agnostic representations compared to BLOOM-7B1 Scao et al. (2022a), which can be attributed to the balanced distribution of our pre-training data. However, for languages with extremely limited resources and poor evaluation performance, such as Bengali and Tamil, FuxiTranyu-8B tends to allocate fewer neurons to process them. Additionally, different layers and components of FuxiTranyu-8B handle multilingual text differently, with deep layers being more language-specific and the importance of attention and MLP components varying across layers.
2 Related Work
The rapid advancement of LLMs has led to a surge in research on multilingual LLMs, aimed at supporting a broader range of languages and tasks. Training multilingual LLMs typically involves a multi-stage process, combining different approaches to enhance the model’s capabilities across multiple languages, either training a model from random initialization on massive multilingual data (e.g., BLOOM Scao et al. (2022a), OPT Zhang et al. (2022), PaLM Chowdhery et al. (2022), LLaMA Touvron et al. (2023a)) or building upon existing pretrained LLMs to reduce computational cost (e.g., X-Gen Vu et al. (2022), FinGPT Luukkonen et al. (2023), Cabrita Larcher et al. (2023), Sabia Almeida et al. (2024)). While these methods have made significant strides in bridging the gap between high- and low-resource languages, challenges still remain.
From-scratch pre-training often struggles with the curse of multilinguality, where adding more languages can lead to performance degradation for low-resource languages. Continual pre-training, while more efficient, suffers from catastrophic forgetting, where models forget previously learned knowledge. Supervised fine-tuning (SFT) often leverages multilingual instruction data or incorporates translation tasks to address data scarcity Shen et al. (2023a); Lai et al. (2023); Wang et al. (2022). However, both methods rely heavily on high-quality, diverse datasets, which are often limited for many languages. Reinforcement Learning from Human Feedback (RLHF) is increasingly used to align models with human preferences Shen et al. (2023b). In multilingual LLMs, multilingual RLHF data are used to train multilingual reward models Chen et al. (2024). However, RLHF typically relies on human-annotated data, which can be expensive and time-consuming to collect, especially for under-resourced languages. Downstream fine-tuning involves either tuning all parameters on downstream tasks Rosenbaum et al. (2022); Yang et al. (2023) or employing parameter-efficient fine-tuning methods to reduce costs Tu et al. (2024); Whitehouse et al. (2023). While these methods can achieve impressive performance, they can also be computationally expensive and may not generalize well to unseen tasks or languages.
Recent years have witnessed that prominent MLLMs have been developed, each with specific training methodologies and strengths. These include BLOOM (176B parameters, open-source, over 46 languages), LLaMA (65B parameters, efficient architecture), PaLM (540B parameters, wide benchmark success), OPT (175B parameters, open-source), Qwen (14B parameters, strong benchmark performance ), Mistral (7B parameters, open-source, competitive performance ), and Orion-14B (14B parameters, diverse data of 2.5T tokens, data scheduling strategy). While these models have achieved impressive results, future work should focus on addressing the limitations of existing approaches. We strongly suggest that efforts should be made to develop more robust and efficient training methods and strategies that address the curse of multilinguality, mitigate catastrophic forgetting, alleviate data imbalance, and minimize reliance on expensive annotated data, especially for low-resource languages.
ISO-931 Language Language Family ISO-931 Language Language Family
ar Arabic Afro-Asiatic ky Kyrgyz Turkic
bg Bulgarian Indo-European lo Lao Kra-Dai
bn Bengali Indo-European ms Malay Austronesian
ca Catalan Indo-European my Burmese Sino-Tibetan
cs Czech Indo-European nl Dutch Indo-European
de German Indo-European pl Polish Indo-European
el Greek Indo-European pt Portuguese Indo-European
en English Indo-European ro Romanian Indo-European
es Spanish Indo-European ru Russian Indo-European
fa Persian Indo-European sv Swedish Indo-European
fi Finnish Uralic ta Tamil Dravidian
fr French Indo-European tg Tajik Indo-European
he Hebrew Afro-Asiatic th Thai Kra-Dai
hi Hindi Indo-European tk Turkmen Turkic
hu Hungarian Indo-European tl Filipino Austronesian
id Indonesia Austronesian tr Turkish Turkic
it Italian Indo-European uk Ukrainian Indo-European
ja Japanese Japanic ur Urdu Indo-European
kk Kazakh Turkic uz Uzbek Turkic
km Khmer Austroasiatic vi Vietnamese Austroasiatic
ko Korean Koreanic zh Chinese Sino-Tibetan
ku Kurdish Indo-European
Table 2: The list of 43 natural languages supported by FuxiTranyu.
Language Size (GB) Ratio (%) Language Size (GB) Ratio (%)
Java 96 17.94 Go 26 4.86
JavaScript 70 13.08 SQL 11 2.06
Python 63 11.77 Rust 9.1 1.70
PHP 59 11.02 Ruby 7.9 1.48
C 53 9.90 Scala 5.1 0.95
C++ 52 9.72 Lua 3.0 0.56
C# 48 8,97 Assembly 1.6 0.30
TypeScript 29 5.42 Visual Basic 1.5 0.28
Table 3: The list of 16 programming languages covered in FuxiTranyu, including the sizes and ratios of each language.
3 Pretraining
We present the strategy we used to determine which languages should be supported by FuxiTranyu series of models in Section 3.1. After that, we elaborate the sources and domains of our pre-training data, and the efforts we have made in the pre-processing stage in Section 3.2. Next, we discuss the details of our tokenizer training in Section 3.3 and the details of our FuxiTranyu architecture in Section 3.4. Finally, we present the pre-training settings in Section 3.5.
3.1 Supported Languages in FuxiTranyu
Our language selection strategy primarily stems from two distinct perspectives: the availability of pre-training data and geographical considerations. We initially approach language selection from the perspective of available pre-training data. Given that the majority of our pre-training data is sourced from web documents, e.g., CulturaX, we determine the languages for pre-training FuxiTranyu based on the statistical information derived from CulturaX. We select the top 21 languages based on the number of available tokens in descending order. Subsequently, we manually incorporate Asian languages, encompassing those from Southeast Asia, West Asia, and Central Asia, resulting in a total of 43 languages. The complete list can be found in Table 2.
In terms of programming languages, we initially consider all 13 languages included in BLOOM Scao et al. (2022a), such as Java, JavaScript, and Python. Additionally, we include three programming languages (SQL, Assembly, and Visual Basic) due to their high popularity, as indicated by the TIOBE index. [^4] https://www.tiobe.com/tiobe-index/ The complete list of programming languages is provided in Table 3.
Figure 1: Languages and domains distribution in the pre-training data of FuxiTranyu.
3.2 Data Collection
The quantity, diversity, and quality of data have proven the most crucial factors determining the performance of a pre-trained base model Hoffmann et al. (2022); Touvron et al. (2023a, b). In pursuit of these objectives, we collect a substantial volume of multilingual data to ensure there are enough tokens for pre-training, in line with scaling laws. Our data collection encompasses a broad spectrum of domains, including public web documents, encyclopedic content, reports, books, scientific articles, and codes. To ensure the quality of the collected corpora, we have employed heuristic quality filters, learned quality filters, and deduplication processes. The composition of the pre-training data mixture is illustrated in Figure 1, and we will delve into the specifics of data collection and pre-processing in the remaining of this section.
A significant portion of our multilingual data comprises web documents, as they provide a vast amount of data for pre-training, akin to other open-sourced LLMs Touvron et al. (2023a); Bai et al. (2023); Cai et al. (2024); Young et al. (2024). We opt to utilize CulturaX Nguyen et al. (2023), a filtered subset of OSCAR Ortiz Su’arez et al. (2020); Suárez et al. (2019) (itself a subset of Common Crawl) and mC4 Raffel et al. (2020) datasets. To enhance the quality and diversity of our pre-training corpora, we further collect data from various sources such as ROOTS Laurençon et al. (2022), MultiUN Eisele and Chen (2010); Chen and Eisele (2012), and OpenSubtitles Lison and Tiedemann (2016). We primarily select documents in languages included in our language list. We further include data sourced from encyclopedias and reports. Inspired by the Phi series models Gunasekar et al. (2023), which leverage high-quality data from textbooks to achieve remarkable performance, we also integrate books and articles data into our final data mixture. Approximately 500GB of articles data have been gathered from Semantic Scholar (S2ORC) Lo et al. (2020), and around 10GB of Chinese books data sourced from Fudan Cbook dataset. [^5] https://github.com/FudanNLPLAB/CBook-150K
Multilingual book data are obtained from Project Gutenberg based on the provided language identity, although it constitutes a small portion of our final corpora. Additionally, we collect 535GB of code data from open-source datasets. The primary source is Starcoderdata, [^6] https://huggingface.co/datasets/bigcode/starcoderdata a subset of the Stack dataset Kocetkov et al. (2022) used to train the StarCoder model Li et al. (2023). We also include a subset of Github code from the RedPajama dataset. [^7] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T
At the filtering stage, we primarily employ three different filtering methods, aligning with previous works Scao et al. (2022a); Almazrouei et al. (2023); Bai et al. (2023); Young et al. (2024). The initial filtering phase incorporates heuristic rules to exclude undesired documents. This involves filtering out documents containing URLs or words listed in blacklists, such as stop words or flagged words. Subsequently, we filter documents based on statistical information, including the ratio or number of repeated n-gram characters or words, as well as the document length. Following this, we apply a learned quality filter method based on specific metrics, such as perplexity. In line with the approach taken in BLOOM Scao et al. (2022a), we utilize KenLM Heafield (2011) to compute the perplexity of the documents and subsequently filter out those surpassing the pre-defined threshold.
Upon completion of the quality filter stage, significant efforts are dedicated to data deduplication, as previous studies have emphasized its importance for LLM performance Lee et al. (2022). We employ fuzzy-match deduplication using the MinHash algorithm. However, due to the memory-intensive nature of deduplication, processing the entire dataset at once on a server with limited memory is unfeasible. Yet, processing only a portion of the data will not achieve complete deduplication. To address this challenge, we apply a strategy of multi-turn micro-deduplication. We first split large documents into multiple chunks and maintain a chunk pool. In each turn, we randomly select chunks from the pool and perform deduplication among these chunks. Once processed, these collected chunks are randomly split into multiple chunks and reintegrated into the chunk pool. This procedure is repeated multiple times until the number of filtered-out documents is less than 1%. In practice, we employ multi-turn deduplication primarily for high-resource languages. For low-resource languages, the entire dataset could fit into memory at once due to the limited amount of pre-training data. In the case of code data, we also utilize the MinHash algorithm for data deduplication. Specifically, we leverage the implementation from the bigcode project. [^8] https://github.com/bigcode-project/bigcode-dataset/blob/main/near_deduplication/minhash_deduplication.py
Figure 2: Fertility test results of the tokenizers for FuxiTranyu, Llama-2, and BLOOM.
3.3 Tokenization
We implement the Byte-level Byte-Pair Encoding (BBPE) algorithm using the Hugging Face tokenizer library. Our tokenizer is initiated from GPT-2’s tokenizer, incorporating both pre-tokenization and post-tokenization processes. Notably, we opt not to split numbers into digits. In line with the approach outlined in BLOOM Scao et al. (2022a), we expand the vocabulary size to 250,680 to accommodate multilingual scenarios, thereby mitigating the risk of over-segmentation in low-resource languages.
For training the tokenizer, we randomly sample 1 million documents for each language from our collected data. It’s worth noting that for languages with a total document count being less than 1 million, we utilize all available documents in the training data for the tokenizer.
Following the approach used in BLOOM, we also evaluate the performance of our tokenizer using the fertility metric. To assess its efficacy, we conduct a comparative analysis with the Llama-2 and BLOOM tokenizers. This evaluation involves computing fertility on the same set of documents across different languages. Results are presented in Figure 2, which indicate that the FuxiTranyu tokenizer is more efficient than the others in most languages. Based on our evaluations and interpretability analysis, we believe that the fertility of the tokenizer positively correlates with the model’s performance on specific languages. In the fertility test, we observe that Bengali (bn), Hindi (hi), and Tamil (ta) exhibit high fertility, indicating lower tokenization efficiency in these languages compared to others. Consequently, the performance and importance of neurons of these languages in our base model are also suboptimal. Further details are discussed in Section 6.1.2.
3.4 Model Architecture
The architecture of FuxiTranyu has been crafted using a modified GPT-2 style framework, drawing inspiration from successful open-source LLMs such as BLOOM, LLaMA, and Qwen. Our modifications are as follows:
  • Untied Embeddings. We opt to separate the weights of the input and output embeddings to enhance performance, despite the resulting increase in total model parameters and memory usage.
  • Linear Bias. In contrast to prior approaches (Chowdhery et al., 2022; Touvron et al., 2023a), we choose not to eliminate the linear bias of the linear projection layers in self-attention and feed-forward layers.
  • Position Encodings. To extend the model’s ability to handle long context, we adopt RoPE Su et al. (2021), replacing the original absolute or relative position embedding method utilized in T5 Raffel et al. (2020). RoPE has demonstrated promising results in managing long context situations and has been widely employed in LLMs Touvron et al. (2023a); Inc. (2023); Bai et al. (2023).
  • Normalization. Given the significance of pre-training stability in training large LMs with a substantial number of tokens, we implement pre-normalization due to its superior stability compared to post-normalization Xiong et al. (2020). Furthermore, we incorporate the widely used RMSNorm Jiang et al. (2023b) to enhance training efficiency.
  • Activation Function. While SwiGLU Shazeer (2020) has been a popular choice for activation functions due to its performance improvements Scao et al. (2022b), it introduces an additional linear function into the activation process, resulting in a 50% increase in parameters in the feed-forward layer. Considering this, we decide to use the GeLU Hendrycks and Gimpel (2016) activation function. GeLU has been shown to achieve similar performance to SwiGLU, as reported in (Scao et al., 2022b).
# Params 8B
Hidden Size 4,096
Intermediate Size 16,384
Heads 32
Layers 30
Position Embed 4,096
Vocab Size 250,752
Learning Rate 3e-4 $\rightarrow$ 1e-4
Batch Size 2M $\rightarrow$ 4M
Context Length 4,096
Training Tokens 606B
FlashAttn V2 $\checkmark$
Table 4: Model size and hyper-parameters. We append 72 dummy tokens to the vocabulary to make the embedding size be divisible by 128.
3.5 Pre-training Details
The training procedure for the FuxiTranyu model adheres to the standard autoregressive language model framework, utilizing the next-token prediction loss as detailed in (Brown et al., 2020). To enhance pre-training efficiency, we employ a document packing method similar to that described in (Raffel et al., 2020). This involves randomly shuffling documents, merging them, and then truncating into multilingual chunks that adhere to a maximum context length of 4096 tokens during the pre-training phase.
To mitigate memory consumption and further improve training efficiency, we leverage ZeRO-2 Rajbhandari et al. (2020) and Flash-Attention V2 Dao (2024) technologies. For optimization, the standard AdamW optimizer Loshchilov and Hutter (2017) is utilized with hyper-parameters set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. We employ the cosine learning rate scheduler, starting with a maximum learning rate of 3e-4 and decaying to a minimum of 10% of the maximum rate. Notably, after encountering divergence issues post-training approximately 241 billion tokens, we reduced the maximum learning rate to 1e-4 to match with the learning rate used in BLOOM, given the multilingual context of both models.
Our FuxiTranyu-8B model is trained using the Megatron-LM Shoeybi et al. (2019) framework on a setup of 32 A800 GPUs, processing a total of 606 billion tokens. The training utilizes FP16 mixed precision to ensure stability. Detailed training parameters and configurations are provided in Table 4.
4 Post-training
To develop a model capable of following instructions and engaging in conversational interactions with humans, we have adopted the instruction fine-tuning and reinforcement learning (RL) approach outlined in (Ouyang et al., 2022).
During the instruction fine-tuning phase, we curate a diverse and high-quality open-source instruction dataset. Given the abundance of instruction-following datasets that have demonstrated exceptional alignment results with various models, manually selecting and fine-tuning the mixture rates for each dataset becomes a challenging task. Consequently, we opt to designate a primary dataset and supplement it with additional datasets. In this context, we select the OpenHermes 2.5 data collection Teknium (2023) as our base dataset, composed of multiple datasets covering a wide range of instructions and yielding excellent results when fine-tuned with Mistral-7B-v0.1. We make modifications to the original OpenHermes 2.5 dataset by replacing Airoboros 2.2 with Airoboros 3.2.999https://huggingface.co/datasets/jondurbin/airoboros-3.2 Additionally, we incorporate the Aya dataset Singh et al. (2024) to enhance the multilingual capabilities of our base model. We filter out the instructions where language is not included in our pre-training language list. To bolster the model’s proficiency in Chinese, we include the COIG-CQIA Bai et al. (2024), ruozhiba-gpt4101010https://huggingface.co/datasets/hfl/ruozhiba_gpt4, and in-house Chinese multidisciplinary instruction data as supplementary datasets. To enhance math and coding abilities, we use the dart-math-hard Tong et al. (2024) and Magicoder-Evol-Instruct  [^11] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110KLuo et al. (2023) datasets.
Models m-ARC m-Hellaswag m-MMLU XWinograd XCOPA XStoryCloze
(25-shot) (10-shot) (5-xhot) (5-shot) (0-shot) (0-shot)
Llama-2-7B 35.5 48.6 35.4 78.0 58.9 55.6
Mistral-7B-v0.1 40.7 54.5 46.7 80.5 55.8 60.2
BLOOM-7B1 31.8 43.4 27.1 70.0 56.9 58.2
PolyLM-13B 30.6 46.0 26.4 73.4 58.9 56.4
LLaMAX2-7B 33.1 50.3 26.7 76.9 54.5 58.8
FuxiTranyu-8B 32.7 51.8 26.6 76.1 60.5 58.9
Table 5: Average performance of FuxiTranyu-8B base model compared to BLOOM-7B1, PolyLM-13B, Llama-2-7B, Mistral-7B-v0.1, and LLaMAX2-7B on mutlilingual discriminative and generative tasks.
In the RL training stage, we opt to use DPO Rafailov et al. (2023) as our RL algorithm instead of RLHF Ouyang et al. (2022); Schulman et al. (2017), as it requires less GPU memory than RLHF, which utilizes PPO as the RL algorithm. We use UltraFeedback Cui et al. (2023) for the DPO training, since this dataset focuses on general alignment ability and has been successfully utilized by Zephyr Tunstall et al. (2023) to train the DPO model.
We leave the settings of post-training in Appendix A.
5 Experiments
Models m-ARC m-Hellaswag m-MMLU XWinograd XCOPA XStoryCloze Translation Summarization
(25-shot) (10-shot) (5-shot) (5-shot) (0-shot) (0-shot) (BLEU, 0-shot) (ROUGE, 0-shot)
Llama-2-Chat-7B 36.4 46.3 36.0 74.8 55.9 56.5 22.1 4.6
Mistral-7B-Instruct-v0.1 36.3 45.5 39.0 74.0 54.5 53.4 19.1 2.2
BLOOMZ-7B1 31.2 38.0 25.8 64.0 53.3 49.8 14.7 4.4
PolyLM-MultiAlpaca-13B 28.6 39.1 25.9 70.9 59.9 57.0 - -
LLaMAX2-Alpaca-7B 38.7 52.5 35.4 77.4 56.6 62.0 29.1 0.3
FuxiTranyu-8B-SFT 31.8 51.5 26.8 75.7 61.3 56.6 25.9 8.9
FuxiTranyu-8B-DPO 32.8 52.2 27.3 74.1 62.1 56.9 26.4 7.3
Table 6: Average performance of FuxiTranyu-8B instruct and chat models compared to BLOOMZ-7B1, Llama-2-Chat-7B, and Mistral-7B-Instruct-v0.1 on mutlilingual discriminative and generative tasks.
We conducted extensive experiments to evaluate the capabilities of FuxiTranyu under the multilingual setting, specifically from the base model to the instruction-tuned model. We selected several models as benchmarks to compare our models with both English-centric and multilingual models. For English-centric models, we compared FuxiTranyu against Llama-2 (Llama-2-7B, Llama-2-chat-7B) (Touvron et al., 2023b) and Mistral (Mistral-7B-v0.1, Mistral-7B-instruct-v0.1) (Jiang et al., 2023a). For multilingual models, we compared FuxiTranyu with BLOOM (BLOOM-7B1, BLOOMZ-7B1) (Scao et al., 2022a; Muennighoff et al., 2022), PolyLM (PolyLM-13B, PolyLM-MultiAlpaca-13B) (Wei et al., 2023), and LLaMAX2 (LLaMAX2-7B, LLaMAX2-7B-Alpaca) (Lu et al., 2024). [^12] LLaMAX series models are continual pre-trained on the Llama-2 model to support beyond 100 languages. We used the LM Evaluation Harness framework (Gao et al., 2023) for all evaluation experiments.
Discriminative Tasks
For evaluating discriminative tasks, we used ARC (Clark et al., 2018), Hellaswag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), XWinograd (Tikhonov and Ryabinin, 2021), XCOPA (Ponti et al., 2020), and XStoryCloze (Lin et al., 2021) datasets. Specifically for the multilingual evaluation, we utilized the multilingual version of ARC, HellaSwag and MMLU datasets (Lai et al., 2023) and selected 15 languages for the evaluation (ar, bn, de, en, es, fr, hu, id, it, pt, ru, sk, ta, vi, zh). For XWinograd, XCOPA, and XStoryCloze datasets, we utilized all of the languages provided in the datasets.
Generative Tasks
We evaluated the performance towards generative tasks, especially in translation and summarization tasks. For translation task, we employed WMT14 in en-fr translation direction (Bojar et al., 2014), WMT16 in en-de and en-ro translation directions (Bojar et al., 2016) and IWSLT 2017 (Cettolo et al., 2017) in en-ar translation direction for measuring the translation performance in our models and benchmark models. For summarization task, we used XL-Sum (Hasan et al., 2021) dataset. We selected 15 languages for the evaluation (ar, en, es, fr, gu, hi, id, mr, pt, ru, sr, ta, uk, vi, zh).
5.1 Base Model Evaluation
First, we report experiment results of our base models vs. baseline models. We focus on evaluating the capabilities of LLMs towards discriminative tasks. Evaluation results are shown in Table 5. Our model achieves the best performance on the XCOPA task. For other tasks, our model is significantly better than multilingual models like BLOOM-7B and PolyLM-13B. When compared to LLaMAX-7B, the evaluation results of our model are almost comparable, with no significant difference from the evaluation results of LLaMAX-7B. But compared with english-centric models, our model still worse than Llama-2-7B and Mistral-7B-v0.1 due to the limited training data used for English.
5.2 Instruction-Tuned Model Evaluation
We further compared our instruction-tuned model with other instruction-tuned models. We evaluated these models on both discriminative and generative tasks. Results are shown in Table 6. On discriminative tasks, our models achieve the best result on XCOPA. For m-Hellaswag, XWinograd, and XStoryCloze, our models outperforms the English-centric models, but slightly underperforms the multilingual models compared with LLaMAX2-7B. Our models still underperforms in m-ARC and m-MMLU tasks due to the limited training data used.
In generative tasks, our model excells on the summarization task, outperforming all baseline models. For the translation task, our model outperforms the English-centric models, but slightly underperforms the multilingual model like LLaMAX2-Alpaca-7B.
More details of our evaluations are discussed in Appendix B, where we report the results for each language tested.
6 Analysis and Interpretability
We further conducted an interpretability analysis of FuxiTranyu to provide a deep understanding of the underlying mechanisms driving its multilingual capabilities. To ensure a comprehensive analysis and consistency with prior research, we investigated our models from both the neuron Wu et al. (2023); Shi et al. (2024); Leng and Xiong (2024); Zhang et al. (2024); Tang et al. (2024); Liu et al. (2024); Kojima et al. (2024) and representation Conneau et al. (2020); Tiyajamorn et al. (2021); Chang et al. (2022); Rajaee and Pilehvar (2022); Xu et al. (2023); Dong et al. (2024); Xie et al. (2024) perspectives. Specifically, our neuron analysis explores the importance of different neurons to multilingual abilities of the model, while the representation analysis examines the characteristics of multilingual representations learned by the model. Here, we first introduce the details and results of our neuron analysis, while the representation analysis is discussed in Section 6.2.
6.1 Neuron Analysis
Neurons in a neural network are the basic computational units of the model. Different inputs may fire neurons in different regions, leading to varied outputs. This computational process can be understood from another perspective: different sets of neurons in the model hold varying degrees of importance for the inputs, thus producing different responses and outputs. To better understand why models generate specific outputs for specific inputs in a multilingual context, we aim to reveal the model’s internal mechanisms by evaluating the importance of neurons. Specifically, we assess the importance of different neurons for various linguistic inputs to determine which neurons play a key role in processing particular languages.
We draw on the approach of assessing parameter sensitivity in model pruning, where the basic idea is that a parameter is considered sensitive or important if removing it, by setting the representation produced by that parameter to zero, significantly affects the loss function Zhang et al. (2024). Specifically, the model can be represented as a parameter set $\bm{\theta}=[\bm{\theta}_{1},\bm{\theta}_{2},\dots,\bm{\theta}_{n}]$, where $\bm{\theta}_{i}\in\mathbb{R}^{d}$ is the $i$-th neuron in the model. Let $\bm{h_{i}}$ denote the representation produced by neuron $\bm{\theta}_{i}$. The importance of neuron $\bm{\theta}_{i}$, denoted as $\Phi(i)$, is defined as the change in the loss function $\mathcal{L}$ before and after setting representation $\textbf{h}_{i}$ to zero. Formally, $\Phi(i)$ can be estimated as follows:
$\Phi(i)=|\Delta{\mathcal{L}(\textbf{h}_{i})}|=\left|\mathcal{L}\left(\textbf{H},\textbf{h}_{i}=\textbf{0}\right)-\mathcal{L}\left(\textbf{H},\textbf{h}_{i}\right)\right|$ (1)
where H is the representation produced by a neuron other than $\bm{\theta}_{i}$ in the same structure as the $\bm{\theta}_{i}$.
Calculating the importance of each neuron in the model using the aforementioned method is very time-consuming, as it requires traversing each neuron. However, based on prior studies, we can simplify these calculations using a Taylor expansion, as shown in Equation 2:
$\begin{split}\Phi(i)&=|\mathcal{L}(\textbf{H},\textbf{h}_{i}=\textbf{0})-(\mathcal{L}(\textbf{H},\textbf{h}_{i}=\textbf{0})\\ &+\frac{\partial\mathcal{L}(\textbf{H},\textbf{h}_{i})}{\partial\textbf{h}_{i}}\textbf{h}_{i}+R_{1}(\textbf{h}_{i}))|\end{split}$ (2)
After ignoring the term $R_{1}(\textbf{h}_{i})$, the neuron importance evaluation function is simplified to $\frac{\partial\mathcal{L}(\textbf{H},\textbf{h}_{i})}{\partial\textbf{h}_{i}}\textbf{h}_{i}$, which is the product of the gradient and the representation. This enables parallel computation of each neuron’s importance.
Furthermore, to measure the significance of a specific parameter set $\bm{\alpha}=[\bm{\theta}_{l},\bm{\theta}_{l+1},\dots,\bm{\theta}_{k}]\subseteq\bm{\theta}$, we compute the importance of each neuron in the set using the following equation:
$\Phi(\bm{\alpha})=\sum_{i=l}^{k}\Phi(i)$ (3)
where $\Phi(\bm{\alpha})$ denotes the importance of the parameter set $\bm{\alpha}$. The set $\bm{\alpha}$ can represent a component or a layer of the model, with the neuron indices in $\bm{\alpha}$ generally being continuous.
6.1.1 Analysis Setup
We chose the Flores-200 dataset (Costa-jussà et al., 2022) to evaluate the importance of neurons. By selecting the languages ar, bn, es, fr, id, pt, ta, vi, zh, en, de, hu, it, ru, and sk, we analyzed the significance of different model components and layers in response to various linguistic inputs.
6.1.2 Results
We analyzed the varying importance of different layers across diverse language inputs, as shown in Figure 4 (Appendix C). Our findings indicate that, universally, shallow layers exhibit low significance while deep layers demonstrate great importance. Notably, languages such as bn and ta exhibit a notably diminished importance in deep layers compared to others, aligning with our evaluation results where these languages perform poorly. This discrepancy may stem from their relatively limited representation learning in the pre-training data.
We then analyzed the significance of various components across different language inputs, depicted in Figure 5 (Appendix C), with 8 components per layer. Our findings mirror previous conclusions: components in shallow layers exhibit low importance, whereas those in deep layers show high significance. Moreover, a more detailed observation reveals that MLP components hold greater importance in shallow layers, whereas attention components are more critical in deep layers.
Figure 3: Similarity distribution of multilingual representations in the intermediate layers of BLOOM-7B1 and FuxiTranyu-8B, with languages sorted based on their percentages in the pre-training data.
6.2 Representation Analysis
Language models encode textual symbols into high-dimensional representations with rich semantic information. For a multilingual language model, due to parameter sharing mechanisms, it encodes textual symbols from different languages into a unified representation space. Furthermore, through multilingual joint training, the model learns multilingual representations, which encode the intrinsic characteristics of languages and the relationships between different languages. Here, we explore the multilingual characteristics of the model from the perspective of the multilingual representations it learns. Specifically, we calculate the similarity of representations across different languages.
To quantitatively evaluate the similarity between different language representations, we choose cosine similarity for its simplicity and effectiveness. To mitigate the impact of semantic differences on our analysis, we collect multilingual text data from open-source parallel corpora. For a language $l$, we input its corresponding text data into the model and collect text representations from the last token of each respective text. We then compute the average of these text representations to obtain the language representation $\bm{v}_{l}$ for language $l$. Finally, we calculate the similarity between two language representations as $\text{sim}(l_{1},l_{2})=\frac{\bm{v}_{1}^{\top}\bm{v}_{2}}{\|\bm{v}_{1}\|\|\bm{v}_{2}\|}$. It’s important to note that we extract language representations and compute similarity across each layer of the model.
6.2.1 Analysis Setup
We selected the Flores-200 dataset (Costa-jussà et al., 2022) as our parallel data source, which includes 2009 sentences for each language. For the explored languages, we chose en, zh, de, fr, es, ru, it, pt, nl, pl, ja, vi, cs, tr, hu, el, sv, ro, uk, and hi, based on their highest language proportions in our pre-training data. For comparison, we also analyzed the BLOOM-7B1 model (Scao et al., 2022a). For this model, we considered en, zh, fr, es, ru, pt, nl, pl, ja, vi, cs, tr, hu, el, sv, ro, uk, hi, fi, and th.
6.2.2 Results
Figure 3 illustrates the similarities distribution of multilingual representations in the intermediate layers of two models, with languages ordered according to the amount of language resources. It is apparent that for the BLOOM-7B, lower multilingual representation similarities tend to occur between the top 10 languages with higher resource availability and the bottom 10 languages with lower resource availability. In contrast, our model learn more consistent multilingual representations for all the languages we explored. This indicates that our model possesses a higher degree of multilingual balance, which is also reflected in our multilingual evaluation results and pre-training corpus.
Furthermore, we calculate the average similarity for each layer of the two models, as shown in Figure 6 (Appendix C). For our model, it can be observed that there is a significant increase in similarity from the embedding layer to layer 0, reaching a very high level. As the depth of the model increases, the similarity continues to rise, indicating that the model learns richer multilingual alignment information in these layers. Subsequently, there is a sharp decrease in similarity from layer 28 to layer 29, suggesting that language-specific multilingual representations in the final layer are learned to predict the diverse multilingual vocabulary. For BLOOM-7B1, the trend of similarity changes across layers is similar, initially increasing and then decreasing, but the changes are more gradual in magnitude.
7 Conclusion
In this paper, we have presented the FuxiTranyu models to address the need for open-source multilingual LLMs. Along with the base model, FuxiTranyu-8B, we also present the fine-tuned models on multilingual supervised fine-tuning dataset and preference dataset, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO. Evaluations on multilingual benchmarks show FuxiTranyu models outperform previous multilingual and monolingual LLMs. Furthermore, interpretability analyses underscore the efficacy of the multilingual capabilities embedded in FuxiTranyu.
Acknowledgements
The present research was supported by the National Key Research and Development Program of China (Grant No. 2023YFE0116400). The computing resources used in this project are supported by the Scientific Computing Center of CIC, Tianjin University.
Appendix A Post-Training Details
During the instruction tuning phase, we executed the fine-tuning process on 5 A100 80GB GPUs, leveraging the TRL framework for instruction fine-tuning and DPO training. Throughout both stages, we employed the ChatML format [^13] https://github.com/openai/openai-python/blob/release-v0.28.0/chatml.md for the chat template, and designated <PAD> as the pad token. We used AdamW Loshchilov and Hutter (2017) optimizer, complemented by a cosine learning rate scheduler. The maximum sequence length was set to 4096 for both stages.
In the SFT stage, we configured the maximum learning rate to 2e-5, with a warmup phase spanning 10% of the total steps. The global batch size was set to 320, and the model was trained for 2 epochs. To optimize memory usage, we enabled Flash-Attention V2 Dao (2024), ZeRO stage 2 Rajbhandari et al. (2020), and gradient checkpointing. Additionally, we employed NEFTune Jain et al. (2023), which introduces noise to embedding weights to enhance the final performance of our instruction-tuned model.
In the subsequent DPO training stage, we adhered to the latest hyper-parameters specified for reproducing the results of Zephyr, as provided by the alignment-handbook.141414alignment_handbook2023 The beta value for DPO was set to 0.01, and the training took 1 epoch on UltraFeedback. The maximum learning rate was set to 5e-7, with a warmup phase covering 10% of the total training steps. Similar to the SFT stage, the global batch size was maintained at 320, and we activated Flash-Attention V2 and gradient checkpointing to optimize memory usage. To accommodate the policy and reference model within memory constraints, we utilized ZeRO stage 3 for the policy model and omitted ZeRO for the reference model.
Appendix B Detailed Evaluation Results
We provide detailed evaluation results for each language in this section. First, we present the results for all 15 tested languages on the multilingual ARC in Table 7, comparing base models and instruction-tuned models. The results show that our models perform better in 1 of the 15 tested languages for the ARC task. We speculate that our models still underforms on this task due to the relatively small amount of training data used.
Models ar bn de en es fr hu id
Base Model
[…]
Mistral-7B-v0.1 30.5 23.4 43.1 60.0 52.5 47.7 38.7 39.0
[…]
PolyLM-13B 27.3 22.4 32.8 41.8 33.2 32.7 23.6 32.8
LLaMAX2-7B 24.4 24.1 35.1 48.7 38.7 38.8 31.6 31.4
[…]
Instuction-tuned Model
Llama-2-Chat-7B 26.2 23.9 39.8 53.6 43.0 42.5 32.4 35.4
[…]
PolyLM-MultiAlpaca-13B 27.4 18.4 30.5 38.2 32.9 32.8 18.6 30.2
[…]
FuxiTranyu-8B-DPO 32.4 26.9 33.8 36.3 35.3 35.5 34.0 33.7
Models it pt ru sk ta vi zh
Base Model
[…]
Mistral-7B-v0.1 49.9 47.2 42.1 37.1 25.9 31.3 42.8
[…]
Instuction-tuned Model
[…]
PolyLM-MultiAlpaca-13B 32.6 32.7 32.5 20.3 20.5 28.8 32.5
[…]
FuxiTranyu-8B-SFT 33.7 33.3 31.1 28.2 23.4 31.9 34.6
FuxiTranyu-8B-DPO 34.6 34.2 32.5 29.3 24.6 32.5 36.9
Table 7: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B, PolyLM-13B, and LLaMAX2-7B models on multilingual ARC (25-shot).
Next, we present the results for all 15 tested languages on multilingual HellaSwag in Table 8, comparing base models and instruction-tuned models. Despite our FuxiTranyu-8B model being trained on only about 600B tokens, it achieves remarkable performance. The SFT and RL-trained models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also deliver promising results across all languages, even competing with powerful monolingual LLMs like Llama-2-7B and Mistral-7B-v0.1, with English language as exception.
Models ar bn de en es fr hu id
Base Model
[…]
PolyLM-13B 39.6 28.4 49.5 71.3 55.8 54.8 29.3 50.1
LLaMAX2-7B 43.3 32.3 53.8 75.4 59.0 58.1 44.1 51.0
[…]
Instuction-tuned Model
Llama-2-Chat-7B 31.4 28.3 50.7 78.6 58.1 57.0 39.0 44.5
Mistral-7B-Instruct-v0.1 31.2 28.7 52.2 70.1 58.1 57.6 39.8 38.1
BLOOMZ-7B1 39.5 31.5 33.1 46.6 48.7 45.7 29.8 42.0
PolyLM-MultiAlpaca-13B 34.0 25.7 40.7 66.0 43.5 43.1 26.7 40.0
[…]
FuxiTranyu-8B-DPO 48.1 33.6 57.7 57.8 62.5 62.5 49.3 54.5
Models it pt ru sk ta vi zh
Base Model
Llama-2-7B 56.0 56.7 49.9 39.2 28.4 45.7 48.7
[…]
PolyLM-13B 51.4 53.7 48.7 30.1 28.0 46.8 52.0
LLaMAX2-7B 56.1 56.8 51.1 47.8 30.0 47.2 49.3
[…]
Instuction-tuned Model
Llama-2-Chat-7B 53.7 54.0 47.6 36.4 28.8 41.2 45.1
Mistral-7B-Instruct-v0.1 54.6 55.8 49.6 37.4 27.7 36.1 45.9
BLOOMZ-7B1 40.3 37.3 33.1 29.6 29.5 40.6 42.6
[…]
Table 8: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B, PolyLM-13B, and LLaMAX2-7B models on multilingual HellaSwag (10-shot).
We report results on multilingual MMLU in Table 9. Our models still underperforms baseline models for all languages. It is in line with the number of training tokens utilized in the pre-training process.
Models ar bn de en es fr hu id
Base Model
[…]
Mistral-7B-v0.1 35.8 32.2 51.7 60.7 53.7 53.5 46.8 46.9
[…]
PolyLM-13B 26.7 26.3 26.1 27.2 26.9 27.2 26.4 24.9
LLaMAX2-7B 25.5 26.2 27.0 28.3 27.0 26.7 26.9 26.8
FuxiTranyu-8B 26.3 25.5 27.6 27.1 27.1 27.5 26.4 26.2
Instuction-tuned Model
[…]
BLOOMZ-7B1 24.4 25.9 25.6 22.7 27.1 27.7 26.1 26.3
PolyLM-MultiAlpaca-13B 25.9 26.6 26.2 25.9 26.5 26.3 25.2 25.4
LLaMAX2-7B-Alpaca 30.0 30.4 36.4 43.0 37.2 36.9 47.6 35.5
FuxiTranyu-8B-SFT 26.0 27.1 26.6 27.0 26.4 27.8 27.3 26.3
FuxiTranyu-8B-DPO 27.0 27.3 27.2 27.0 27.4 27.8 27.6 26.4
Models it pt ru sk ta vi zh
Base Model
[…]
Mistral-7B-v0.1 52.7 53.4 49.8 45.4 29.7 41.5 46.0
BLOOM-7B1 25.7 25.3 26.2 26.1 26.6 28.1 29.1
PolyLM-13B 27.5 24.5 26.3 27.4 26.4 25.3 26.8
LLaMAX2-7B 27.0 26.9 27.0 26.6 26.2 26.8 26.1
FuxiTranyu-8B 27.1 26.8 27.7 26.0 26.3 26.3 26.0
Instuction-tuned Model
[…]
BLOOMZ-7B1 25.8 22.8 25.4 26.3 26.7 26.3 27.2
PolyLM-MultiAlpaca-13B 25.9 26.2 26.2 25.5 25.5 25.7 26.1
[…]
FuxiTranyu-8B-SFT 27.1 27.0 26.8 27.2 26.4 25.9 27.0
FuxiTranyu-8B-DPO 27.5 27.7 28.0 27.6 26.9 26.2 27.7
Table 9: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B, PolyLM-13B, and LLaMAX2-7B models on multilingual MMLU (5-shot).
Results on XWinograd are depicted in Table 10. Our FuxiTranyu SFT and DPO models achieve better results in Portuguese and Chinese. Although our models underperforms in English, French, Russian, and Japanese compared to Llama-2-7B, they outperforms previous multilingual LLMs like BLOOM-7B1 and PolyLM-13B across all languages.
Models fr pt zh en ru jp
Base
[…]
Mistral-7B-v0.1 81.9 80.6 80.0 90.6 72.4 77.5
BLOOM-7B1 71.1 76.8 74.4 82.2 56.8 58.5
PolyLM-13B 73.5 74.9 76.6 84.6 65.1 65.7
[…]
Instuction-tuned Model
[…]
BLOOMZ-7B1 68.7 65.4 71.0 83.5 53.7 56.4
PolyLM-MultiAlpaca-13B 71.1 72.2 73.6 83.9 67.9 65.2
[…]
Table 10: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XWinograd (5-shot).
Results on XCOPA and XStoryCloze are shown in Table 11 and Table 12. For XCOPA, our base models achieve better results in sw, ta, tr, and vi. When compared to instruction-tuned models, our models achieve better results in more languages, specifically in it, id, ta, th, tr, vi, and zh. On the XStoryCloze task, our base models achieve better results in three languages: ar, my, and ru. However, for instruction-tuned models, our models outperforms other baseline models only in my.
Models et ht it id qu sw ta th tr vi zh
Base
[…]
Instuction-tuned Model
Llama-2-Chat-7B 47.8 51.4 67.0 62.4 50.8 52.2 50.6 54.8 55.6 61.6 61.2
Mistral-7B-Instruct-v0.1 48.2 51.2 65.4 54.0 49.2 54.6 55.2 53.2 52.2 53.2 63.4
[…]
LLaMAX-7B-Alpaca 51.2 54.2 61.0 57.2 52.4 55.0 57.0 56.4 55.4 55.4 67.6
[…]
Table 11: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XCOPA (0-shot).
Models ar es eu hi id my ru sw te zh
Base
[…]
Instuction-tuned Model
[…]
Mistral-7B-Instruct-v0.1 47.1 63.3 50.0 49.8 52.3 47.6 62.3 49.6 51.8 59.7
[…]
Table 12: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XStoryCloze (0-shot).
We present our evaluation results for generative tasks in Table 13 and Table 14. On the XL-Sum task, our models significantly outperform all baseline models across all evaluated languages, demonstrating the potential of our models for summarization task, particularly in a multilingual context. For the translation tasks in WMT14, WMT16, and IWSLT2017, our models excell in the en-ro, en-de, and en-fr translation directions. However, they still lag behind other baseline models in the ro-en, de-en, fr-en, ar-en, and en-ar translation directions. This indicates that our models perform significantly better for out-of-English translation directions. Although our models underperform in the en-ar direction compared to LLaMAX-2-Alpaca, they still achieve notably better results than other models.
Models ar en es fr gu hi id mr pt ru sr ta uk vi zh
[…]
Mistral-7B-Instruct-v0.1 0.1 11.0 3.0 3.4 0.3 0.2 3.1 0.6 3.2 0.4 2.1 0.2 0.3 4.6 0.6
[…]
LLaMAX2-7B-Alpaca 0.0 1.7 0.5 0.7 0.0 0.0 0.3 0.0 0.2 0.0 0.5 0.1 0.1 0.2 0.0
[…]
Table 13: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XL-Sum (0-shot).
Models WMT16 (EN-RO) WMT16 (RO-EN) WMT16 (EN-DE) WMT16 (DE-EN)
BLEU CHRF BLEU CHRF BLEU CHRF BLEU CHRF
[…]
Mistral-7B-Instruct-v0.1 13.66 41.47 24.58 53.04 19.41 49.25 30.19 58.27
BLOOMZ-7B1 1.88 20.09 11.35 36.22 3.76 23.27 22.30 46.69
LLaMAX2-7B-Alpaca 24.52 51.94 36.02 60.85 26.31 53.95 37.05 61.90
[…]
Models WMT14 (EN-FR) WMT14 (FR-EN) IWSLT2017-AR-EN IWSLT2017-EN-AR
BLEU CHRF BLEU CHRF BLEU CHRF BLEU CHRF
[…]
Mistral-7B-Instruct-v0.1 24.24 52.08 31.40 59.50 9.13 32.64 0.31 13.31
[…]
LLaMAX2-7B-Alpaca 32.86 59.53 36.00 61.64 29.76 52.68 10.47 40.27
FuxiTranyu-8B-SFT 34.06 60.74 28.83 57.86 21.42 42.91 8.19 35.67
[…]
Table 14: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on WMT14, WMT16, and IWSLT2017 (0-shot).
Appendix C Detailed Analysis Results
We present the varying importance of different layers across diverse language inputs in Figure 4. Figure 5 shows the significance of various components across different language inputs, with 8 components per layer. Furthermore, we calculate the average similarity of multilingual representations across model layers, as shown in Figure 6.
Figure 4: Importance of model layers across various language settings.
Figure 5: Importance of model components across various language settings.
Figure 6: Averaged similarity distribution of multilingual representations for each layer of BLOOM-7B1 and FuxiTranyu-8B, with “emb” denoting the embedding layer.

### v3

@@
 To mitigate this challenge, we present FuxiTranyu, an open-source multilingual LLM, which is designed to satisfy the need of the research community for balanced and high-performing multilingual capabilities.
-FuxiTranyu-8B, the base model with 8 billion parameters, is trained from scratch on a meticulously balanced multilingual data repository that contains 600 billion tokens covering 43 natural languages and 16 programming languages.
-In addition to the base model, we also develop two instruction-tuned models: FuxiTranyu-8B-SFT that is fine-tuned on a diverse multilingual instruction dataset, and FuxiTranyu-8B-DPO that is further refined with DPO on a preference dataset for enhanced alignment ability.
-Extensive experiments on a wide range of multilingual benchmarks demonstrate the competitive performance of FuxiTranyu against existing multilingual LLMs, e.g., BLOOM-7B, PolyLM-13B, Llama-2-Chat-7B and Mistral-7B-Instruct.
-Interpretability analyses at both the neuron and representation level suggest that FuxiTranyu is able to learn consistent multilingual representations across different languages.
-To promote further research into multilingual LLMs and their working mechanisms, we release both the base and instruction-tuned FuxiTranyu models together with 58 pretraining checkpoints at HuggingFace [^1] https://huggingface.co/TJUNLP/FuxiTranyu-8B and Github.
+The base model, FuxiTranyu-8B, features 8 billion parameters and is trained from scratch on meticulously balanced multilingual data that contains 600 billion tokens covering 43 natural languages and 16 programming languages.
+We also develop two instruction-tuned models: FuxiTranyu-8B-SFT which is fine-tuned on a diverse multilingual instruction dataset, and FuxiTranyu-8B-DPO which is further refined with DPO on a preference dataset for enhanced alignment ability.
+Extensive experiments on a wide range of multilingual benchmarks demonstrate the competitive performance of FuxiTranyu against existing multilingual LLMs, e.g., BLOOM-7B, PolyLM-13B, and Mistral-7B-Instruct.
+Both neuron and representation interpretability analyses reveal that FuxiTranyu achieves consistent multilingual representations across languages.
+To promote further research into multilingual LLMs, we release both the base and instruction-tuned FuxiTranyu models together with 58 pre-training checkpoints at HuggingFace [^1] https://huggingface.co/TJUNLP/FuxiTranyu-8B and Github.
 [^2] https://github.com/tjunlp-lab/FuxiTranyu
-11footnotetext: Correspondence to: Deyi Xiong.
+11footnotetext: Corresponding author.
 1 Introduction
-A well-pretrained base model plays a pivotal role in facilitating research and applications of large language models.
+A well-pretrained base model is crucial for facilitating research and applications of large language models.
 However, training a base LLM from scratch typically demands a substantial amount of data and significant computational resources, posing a barrier to the development of new LLMs.
-On the other hand, the majority of LLMs are usually tailored to specific languages such as English Touvron et al.
+The majority of LLMs are usually tailored to specific languages such as English Touvron et al.
 (2023a, b) or Chinese Bai et al.
-(2023), neglecting the high demand for multilingual capabilities across multiple languages, especially low-resource languages.
-While certain LLMs, such as Mistral models Jiang et al.
-(2023a), demonstrate multilingual capabilities, their language coverage remains limited.
-This limitation significantly restricts the exploration of multilingualism in LLMs under the massive multilingual setting.
+(2023), neglecting the growing demand for multilingual capabilities, especially from low-resource languages.
+While certain LLMs like Mistral models Jiang et al.
+(2023a) demonstrate multilingual capabilities, their coverage is limited, restricting the exploration in massively multilingual settings.
 LLMs Pre-training Tokens Languages
@@
 Checkpoints Available
-BLOOM-7B1 Scao et al. (2022a) 300B 46 NLs + 13 PLs $\times$
+BLOOM-7B1 Scao et al. (2022a) 300B 46 NLs + 13 PLs
 Aya 23-8B Aryabumi et al. (2024) Unknown 23 NLs $\times$ $\times$
@@
 However, as highlighted by the alignment hypothesis in LIMA Zhou et al.
-(2024), the knowledge and capabilities of LLMs are predominantly derived from pre-training rather than supervised fine-tuning.
-Supervised fine-tuning primarily serves to align the behaviors of these models with instructions, which constitutes a sub-distribution of the pre-training data.
-Consequently, for LLMs whose pre-training data are dominated by a few languages, the effectiveness of supervised fine-tuning in enhancing their multilingual capabilities might be limited.
+(2024), the knowledge of LLMs is predominantly derived from pre-training, while supervised fine-tuning primarily aligns model behavior to instructions, which is a narrow subset of the pre-training data.
+This makes fine-tuning less effective for boosting multilingual abilities when pre-training is dominated by a few languages.
 Other initiatives have focused on pre-training multilingual LLMs, such as BLOOM Scao et al.
@@
 Nevertheless, these efforts are hindered by their performance, which does not measure up to that of current trending LLMs.
-BLOOM suffers from outdated training data while PolyLM is undermined by imbalanced language distribution, with English data accounting for approximately 70% and Chinese for ~20%, potentially leading to insufficient learning of under-represented languages.
+BLOOM suffers from outdated training data, while PolyLM is undermined by imbalanced language distribution, with English data accounting for approximately 70% and Chinese for ~20%, potentially leading to insufficient learning of under-represented languages.
 Previous studies Xu et al.
-(2024) disclose three traits of multilingual LLMs caused by imbalanced language resources in pre-training: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, suggesting that multilingual LLMs could benefit from balanced data distribution across languages.
+(2024) disclose three traits of multilingual LLMs caused by imbalanced language resources: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional transfer between high- and low-resource languages, emphasizing the need for balanced data distribution.
 Recently introduced multilingual LLMs, e.g., Aya 23 models Aryabumi et al.
@@
 Table 1 compares FuxiTranyu with currently available multilingual LLMs from different perspectives.
-In addition to the base model, we develop two instruction-tuned models, FuxiTranyu-8B-SFT that is fine-tuned on a collected high-quality multilingual instruction dataset, and FuxiTranyu-8B-DPO that is further tuned on preferences with DPO for enhanced alignment ability.
-To evaluate multilingual capabilities of the FuxiTranyu models, we have conducted extensive evaluations across multiple domains, encompassing multilingual discriminative tasks such as multilingual ARC, HellaSwag, and MMLU Lai et al.
+In addition to the base model, we develop two instruction-tuned models: FuxiTranyu-8B-SFT, fine-tuned on a collected high-quality multilingual instruction dataset, and FuxiTranyu-8B-DPO, further tuned on preferences with DPO for enhanced alignment ability.
+Our evaluations focus on knowledge, capability and alignment dimensions categorized by Guo et al.
+(2023).
+Evaluation results on multilingual discriminative tasks such as multilingual ARC, HellaSwag, and MMLU Lai et al.
 (2023), XWinograd Muennighoff et al.
@@
 (2017) and XL-Sum summarization benchmark Hasan et al.
-(2021).
-Our evaluations focus on knowledge, capability and alignment dimensions categorized by Guo et al.
-(2023).
-As detailed in Section 5, FuxiTranyu models have demonstrated superior performance on the multilingual ARC, HellaSwag, MMLU, XWinograd, XCOPA, and XStoryCloze compared to BLOOM-7B1 and PolyLM-13B.
-Furthermore, our two instruction-tuned models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, outperform Llama-2-Chat-7B, Mistral-7B-Instruct-v0.1, BLOOMZ-7B1, PolyLM-MultiAlpaca-13B on translation benchmarks.
-FuxiTranyu also achieves remarkable results on summarization.
-To provide a deep understanding of the multilingual capabilities of FuxiTranyu models, we have conducted interpretability analyses from two distinct perspectives: neuron analysis and representation analysis, as detailed in Section 6.
-Analysis results indicate that FuxiTranyu-8B has learned more language-agnostic representations compared to BLOOM-7B1 Scao et al.
-(2022a), which can be attributed to the balanced distribution of our pre-training data.
-However, for languages with extremely limited resources and poor evaluation performance, such as Bengali and Tamil, FuxiTranyu-8B tends to allocate fewer neurons to process them.
+(2021), demonstrate FuxiTranyu’s superior performance compared to BLOOM-7B1 and PolyLM-13B, as detailed in Section 5.
+The instruction-tuned models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also outperform Llama-2-Chat-7B, Mistral-7B-Instruct-v0.1, BLOOMZ-7B1, PolyLM-MultiAlpaca-13B on translation and summarization benchmarks.
+To further understand the multilingual capabilities of FuxiTranyu models, we have conducted neuron- and representation-level analysis, revealing that FuxiTranyu-8B learns more language-agnostic representations compared to BLOOM-7B1 Scao et al.
+(2022a), which can be attributed to the balanced pre-training data.
+However, languages with extremely limited resources, such as Bengali and Tamil, are allocated with fewer neurons.
 Additionally, different layers and components of FuxiTranyu-8B handle multilingual text differently, with deep layers being more language-specific and the importance of attention and MLP components varying across layers.
 2 Related Work
-The rapid advancement of LLMs has led to a surge in research on multilingual LLMs, aimed at supporting a broader range of languages and tasks.
-Training multilingual LLMs typically involves a multi-stage process, combining different approaches to enhance the model’s capabilities across multiple languages, either training a model from random initialization on massive multilingual data (e.g., BLOOM Scao et al.
-(2022a), OPT Zhang et al.
-(2022), PaLM Chowdhery et al.
-(2022), LLaMA Touvron et al.
-(2023a)) or building upon existing pretrained LLMs to reduce computational cost (e.g., X-Gen Vu et al.
-(2022), FinGPT Luukkonen et al.
-(2023), Cabrita Larcher et al.
-(2023), Sabia Almeida et al.
-(2024)).
-While these methods have made significant strides in bridging the gap between high- and low-resource languages, challenges still remain.
-From-scratch pre-training often struggles with the curse of multilinguality, where adding more languages can lead to performance degradation for low-resource languages.
-Continual pre-training, while more efficient, suffers from catastrophic forgetting, where models forget previously learned knowledge.
+Recent advanced LLMs Touvron et al.
+(2023a, b); Dubey et al.
+(2024); Bai et al.
+(2023); Yang et al.
+(2024); Young et al.
+(2024); Jiang et al.
+(2023a); Team et al.
+(2024a, b) have excelled in NLP and cross-modal tasks, sparking increased research on multilingual LLMs Scao et al.
+(2022a); Chowdhery et al.
+(2022); Wei et al.
+(2023), which aim at broader language support.
+There are three main approaches to building multilingual LLMs: pre-training from scratch, continual pre-training, and post-training (e.g., supervised fine-tuning and reinforcement learning from human feedback).
+Pre-training from scratch, like PaLM 1&2 Chowdhery et al.
+(2022); Anil et al.
+(2023), BLOOM Scao et al.
+(2022a), and PolyLM Wei et al.
+(2023), leverages extensive training corpora from diverse sources, enabling the incorporation of new knowledge.
+However, pre-training poses a variety of challenges, such as the need for vast computing resources, which can hinder the development of new multilingual LLMs.
+Additionally, it also suffers from the curse of multilinguality Conneau et al.
+(2019); Chai et al.
+(2022); Dubey et al.
+(2024); Gurgurov et al.
+(2024), where the performance of individual languages deteriorates as the number of languages increases.
+On the other hand, continual pre-training, like Aurora-M Nakamura et al.
+(2024), LLaMAX Lu et al.
+(2024), is more efficient but risks catastrophic forgetting of previously learned knowledge.
 Supervised fine-tuning (SFT) often leverages multilingual instruction data or incorporates translation tasks to address data scarcity Shen et al.
@@
 (2022).
-However, both methods rely heavily on high-quality, diverse datasets, which are often limited for many languages.
+However, both continual pre-training and SFT rely heavily on high-quality, diverse datasets, which are often limited to many languages.
 Reinforcement Learning from Human Feedback (RLHF) is increasingly used to align models with human preferences Shen et al.
@@
 However, RLHF typically relies on human-annotated data, which can be expensive and time-consuming to collect, especially for under-resourced languages.
-Downstream fine-tuning involves either tuning all parameters on downstream tasks Rosenbaum et al.
-(2022); Yang et al.
-(2023) or employing parameter-efficient fine-tuning methods to reduce costs Tu et al.
-(2024); Whitehouse et al.
-(2023).
-While these methods can achieve impressive performance, they can also be computationally expensive and may not generalize well to unseen tasks or languages.
-Recent years have witnessed that prominent MLLMs have been developed, each with specific training methodologies and strengths.
-These include BLOOM (176B parameters, open-source, over 46 languages), LLaMA (65B parameters, efficient architecture), PaLM (540B parameters, wide benchmark success), OPT (175B parameters, open-source), Qwen (14B parameters, strong benchmark performance ), Mistral (7B parameters, open-source, competitive performance ), and Orion-14B (14B parameters, diverse data of 2.5T tokens, data scheduling strategy).
-While these models have achieved impressive results, future work should focus on addressing the limitations of existing approaches.
-We strongly suggest that efforts should be made to develop more robust and efficient training methods and strategies that address the curse of multilinguality, mitigate catastrophic forgetting, alleviate data imbalance, and minimize reliance on expensive annotated data, especially for low-resource languages.
-ISO-931 Language Language Family ISO-931 Language Language Family
-ar Arabic Afro-Asiatic ky Kyrgyz Turkic
-bg Bulgarian Indo-European lo Lao Kra-Dai
-bn Bengali Indo-European ms Malay Austronesian
-ca Catalan Indo-European my Burmese Sino-Tibetan
-cs Czech Indo-European nl Dutch Indo-European
-de German Indo-European pl Polish Indo-European
-el Greek Indo-European pt Portuguese Indo-European
-en English Indo-European ro Romanian Indo-European
-es Spanish Indo-European ru Russian Indo-European
-fa Persian Indo-European sv Swedish Indo-European
-fi Finnish Uralic ta Tamil Dravidian
-fr French Indo-European tg Tajik Indo-European
-he Hebrew Afro-Asiatic th Thai Kra-Dai
-hi Hindi Indo-European tk Turkmen Turkic
-hu Hungarian Indo-European tl Filipino Austronesian
-id Indonesia Austronesian tr Turkish Turkic
-it Italian Indo-European uk Ukrainian Indo-European
-ja Japanese Japanic ur Urdu Indo-European
-kk Kazakh Turkic uz Uzbek Turkic
-km Khmer Austroasiatic vi Vietnamese Austroasiatic
-ko Korean Koreanic zh Chinese Sino-Tibetan
-ku Kurdish Indo-European
-Table 2: The list of 43 natural languages supported by FuxiTranyu.
-Language Size (GB) Ratio (%) Language Size (GB) Ratio (%)
-Java 96 17.94 Go 26 4.86
-JavaScript 70 13.08 SQL 11 2.06
-Python 63 11.77 Rust 9.1 1.70
-PHP 59 11.02 Ruby 7.9 1.48
-C 53 9.90 Scala 5.1 0.95
-C++ 52 9.72 Lua 3.0 0.56
-C# 48 8,97 Assembly 1.6 0.30
-TypeScript 29 5.42 Visual Basic 1.5 0.28
-Table 3: The list of 16 programming languages covered in FuxiTranyu, including the sizes and ratios of each language.
+While these methods can achieve impressive performance, they can also be computationally expensive and may not generalize well to unseen languages.
 3 Pretraining
-We present the strategy we used to determine which languages should be supported by FuxiTranyu series of models in Section 3.1.
-After that, we elaborate the sources and domains of our pre-training data, and the efforts we have made in the pre-processing stage in Section 3.2.
-Next, we discuss the details of our tokenizer training in Section 3.3 and the details of our FuxiTranyu architecture in Section 3.4.
-Finally, we present the pre-training settings in Section 3.5.
-3.1 Supported Languages in FuxiTranyu
-Our language selection strategy primarily stems from two distinct perspectives: the availability of pre-training data and geographical considerations.
-We initially approach language selection from the perspective of available pre-training data.
-Given that the majority of our pre-training data is sourced from web documents, e.g., CulturaX, we determine the languages for pre-training FuxiTranyu based on the statistical information derived from CulturaX.
-We select the top 21 languages based on the number of available tokens in descending order.
-Subsequently, we manually incorporate Asian languages, encompassing those from Southeast Asia, West Asia, and Central Asia, resulting in a total of 43 languages.
-The complete list can be found in Table 2.
-In terms of programming languages, we initially consider all 13 languages included in BLOOM Scao et al.
-(2022a), such as Java, JavaScript, and Python.
-Additionally, we include three programming languages (SQL, Assembly, and Visual Basic) due to their high popularity, as indicated by the TIOBE index.
-[^4] https://www.tiobe.com/tiobe-index/ The complete list of programming languages is provided in Table 3.
+We elaborate on the sources and domains of our pre-training data and the efforts we have made in the pre-processing stage in Section 3.1.
+Next, we discuss the details of our FuxiTranyu architecture in Section 3.2.
+We present the strategy we used to determine which languages should be supported by the FuxiTranyu series of models in Appendix A, the details of our tokenizer training in Appendix B, and the pre-training settings in Section C.1.
 Figure 1: Languages and domains distribution in the pre-training data of FuxiTranyu.
@@
 The composition of the pre-training data mixture is illustrated in Figure 1, and we will delve into the specifics of data collection and pre-processing in the remaining of this section.
-A significant portion of our multilingual data comprises web documents, as they provide a vast amount of data for pre-training, akin to other open-sourced LLMs Touvron et al.
+A significant portion of our multilingual data comprises web documents, a common approach in open-sourced LLMs Touvron et al.
 (2023a); Bai et al.
@@
 (2020) datasets.
-To enhance the quality and diversity of our pre-training corpora, we further collect data from various sources such as ROOTS Laurençon et al.
-(2022), MultiUN Eisele and Chen (2010); Chen and Eisele (2012), and OpenSubtitles Lison and Tiedemann (2016).
-We primarily select documents in languages included in our language list.
-We further include data sourced from encyclopedias and reports.
-Inspired by the Phi series models Gunasekar et al.
-(2023), which leverage high-quality data from textbooks to achieve remarkable performance, we also integrate books and articles data into our final data mixture.
-Approximately 500GB of articles data have been gathered from Semantic Scholar (S2ORC) Lo et al.
-(2020), and around 10GB of Chinese books data sourced from Fudan Cbook dataset.
-[^5] https://github.com/FudanNLPLAB/CBook-150K
-Multilingual book data are obtained from Project Gutenberg based on the provided language identity, although it constitutes a small portion of our final corpora.
-Additionally, we collect 535GB of code data from open-source datasets.
-The primary source is Starcoderdata, [^6] https://huggingface.co/datasets/bigcode/starcoderdata a subset of the Stack dataset Kocetkov et al.
+To improve quality and diversity, we supplement these with data from ROOTS Laurençon et al.
+(2022), MultiUN Eisele and Chen (2010); Chen and Eisele (2012), and OpenSubtitles Lison and Tiedemann (2016), focusing on languages in our language list.
+Additionally, we incorporate data from encyclopedias, reports, books, and articles, drawing inspiration from Phi series models Gunasekar et al.
+(2023) that achieve strong results using high-quality textbooks.
+We have collected approximately 500GB of article data from Semantic Scholar (S2ORC) Lo et al.
+(2020), and around 10GB of Chinese books from the Fudan Cbook dataset.
+[^4] https://github.com/FudanNLPLAB/CBook-150K We also source multilingual book data from Project Gutenberg, though it forms a small portion of the final corpus.
+Additionally, we collect 535GB of code data from open-source datasets, primarily from Starcoder data, [^5] https://huggingface.co/datasets/bigcode/starcoderdata a subset of the Stack dataset Kocetkov et al.
 (2022) used to train the StarCoder model Li et al.
@@
 [^6] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T
-At the filtering stage, we primarily employ three different filtering methods, aligning with previous works Scao et al.
+At the filtering stage, we primarily employ three methods similar to prior works Scao et al.
 (2022a); Almazrouei et al.
@@
 (2024).
-The initial filtering phase incorporates heuristic rules to exclude undesired documents.
-This involves filtering out documents containing URLs or words listed in blacklists, such as stop words or flagged words.
-Subsequently, we filter documents based on statistical information, including the ratio or number of repeated n-gram characters or words, as well as the document length.
+The initial filtering phase uses heuristic rules to exclude undesired documents.
+This involves filtering out documents containing blacklisted URLs or words, such as stop words or flagged words.
+Subsequently, we filter documents based on statistical information, including the ratio or the number of repeated n-gram characters or words, as well as the document length.
 Following this, we apply a learned quality filter method based on specific metrics, such as perplexity.
 In line with the approach taken in BLOOM Scao et al.
-(2022a), we utilize KenLM Heafield (2011) to compute the perplexity of the documents and subsequently filter out those surpassing the pre-defined threshold.
+(2022a), we utilize KenLM Heafield (2011) to compute the perplexity of the documents and subsequently filter out those exceeding a predefined threshold.
 Upon completion of the quality filter stage, significant efforts are dedicated to data deduplication, as previous studies have emphasized its importance for LLM performance Lee et al.
 (2022).
-We employ fuzzy-match deduplication using the MinHash algorithm.
+We employ MinHash for fuzzy-match deduplication.
 However, due to the memory-intensive nature of deduplication, processing the entire dataset at once on a server with limited memory is unfeasible.
@@
 To address this challenge, we apply a strategy of multi-turn micro-deduplication.
-We first split large documents into multiple chunks and maintain a chunk pool.
-In each turn, we randomly select chunks from the pool and perform deduplication among these chunks.
-Once processed, these collected chunks are randomly split into multiple chunks and reintegrated into the chunk pool.
-This procedure is repeated multiple times until the number of filtered-out documents is less than 1%.
-In practice, we employ multi-turn deduplication primarily for high-resource languages.
-For low-resource languages, the entire dataset could fit into memory at once due to the limited amount of pre-training data.
+We split large documents into chunks and store them in a chunk pool.
+In each turn, we randomly select chunks from the pool, assemble them back into documents, and perform deduplication on these assembled documents.
+After processing, the deduplicated documents are again split into chunks and reintegrated into the pool.
+This process is repeated multiple times until the number of filtered-out documents drops below 1%.
+This approach is used for high-resource languages, while low-resource languages are processed in memory due to their smaller dataset size.
 In the case of code data, we also utilize the MinHash algorithm for data deduplication.
@@
 [^7] https://github.com/bigcode-project/bigcode-dataset/blob/main/near_deduplication/minhash_deduplication.py
-Figure 2: Fertility test results of the tokenizers for FuxiTranyu, Llama-2, and BLOOM.
-3.3 Tokenization
-We implement the Byte-level Byte-Pair Encoding (BBPE) algorithm using the Hugging Face tokenizer library.
-Our tokenizer is initiated from GPT-2’s tokenizer, incorporating both pre-tokenization and post-tokenization processes.
-Notably, we opt not to split numbers into digits.
-In line with the approach outlined in BLOOM Scao et al.
-(2022a), we expand the vocabulary size to 250,680 to accommodate multilingual scenarios, thereby mitigating the risk of over-segmentation in low-resource languages.
-For training the tokenizer, we randomly sample 1 million documents for each language from our collected data.
-It’s worth noting that for languages with a total document count being less than 1 million, we utilize all available documents in the training data for the tokenizer.
-Following the approach used in BLOOM, we also evaluate the performance of our tokenizer using the fertility metric.
-To assess its efficacy, we conduct a comparative analysis with the Llama-2 and BLOOM tokenizers.
-This evaluation involves computing fertility on the same set of documents across different languages.
-Results are presented in Figure 2, which indicate that the FuxiTranyu tokenizer is more efficient than the others in most languages.
-Based on our evaluations and interpretability analysis, we believe that the fertility of the tokenizer positively correlates with the model’s performance on specific languages.
-In the fertility test, we observe that Bengali (bn), Hindi (hi), and Tamil (ta) exhibit high fertility, indicating lower tokenization efficiency in these languages compared to others.
-Consequently, the performance and importance of neurons of these languages in our base model are also suboptimal.
-Further details are discussed in Section 6.1.2.
 3.2 Model Architecture
@@
 (2020).
-Furthermore, we incorporate the widely used RMSNorm Jiang et al.
+Furthermore, we incorporate the widely used RMSNorm Zhang and Sennrich (2019); Jiang et al.
 (2023b) to enhance training efficiency.
@@
 GeLU has been shown to achieve similar performance to SwiGLU, as reported in (Scao et al., 2022b).
-# Params 8B
-Hidden Size 4,096
-Intermediate Size 16,384
-Heads 32
-Layers 30
-Position Embed 4,096
-Vocab Size 250,752
-Learning Rate 3e-4 $\rightarrow$ 1e-4
-Batch Size 2M $\rightarrow$ 4M
-Context Length 4,096
-Training Tokens 606B
-FlashAttn V2 $\checkmark$
-Table 4: Model size and hyper-parameters. We append 72 dummy tokens to the vocabulary to make the embedding size be divisible by 128.
-3.5 Pre-training Details
-The training procedure for the FuxiTranyu model adheres to the standard autoregressive language model framework, utilizing the next-token prediction loss as detailed in (Brown et al., 2020).
-To enhance pre-training efficiency, we employ a document packing method similar to that described in (Raffel et al., 2020).
-This involves randomly shuffling documents, merging them, and then truncating into multilingual chunks that adhere to a maximum context length of 4096 tokens during the pre-training phase.
-To mitigate memory consumption and further improve training efficiency, we leverage ZeRO-2 Rajbhandari et al.
-(2020) and Flash-Attention V2 Dao (2024) technologies.
-For optimization, the standard AdamW optimizer Loshchilov and Hutter (2017) is utilized with hyper-parameters set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$.
-We employ the cosine learning rate scheduler, starting with a maximum learning rate of 3e-4 and decaying to a minimum of 10% of the maximum rate.
-Notably, after encountering divergence issues post-training approximately 241 billion tokens, we reduced the maximum learning rate to 1e-4 to match with the learning rate used in BLOOM, given the multilingual context of both models.
-Our FuxiTranyu-8B model is trained using the Megatron-LM Shoeybi et al.
-(2019) framework on a setup of 32 A800 GPUs, processing a total of 606 billion tokens.
-The training utilizes FP16 mixed precision to ensure stability.
-Detailed training parameters and configurations are provided in Table 4.
 4 Post-training
-To develop a model capable of following instructions and engaging in conversational interactions with humans, we have adopted the instruction fine-tuning and reinforcement learning (RL) approach outlined in (Ouyang et al., 2022).
+To develop a model capable of following instructions and engaging in conversational interactions with humans, we have adopted the instruction fine-tuning and reinforcement learning (RL) approach outlined in Ouyang et al.
+(2022).
 During the instruction fine-tuning phase, we curate a diverse and high-quality open-source instruction dataset.
@@
 Consequently, we opt to designate a primary dataset and supplement it with additional datasets.
-In this context, we select the OpenHermes 2.5 data collection Teknium (2023) as our base dataset, composed of multiple datasets covering a wide range of instructions and yielding excellent results when fine-tuned with Mistral-7B-v0.1.
-We make modifications to the original OpenHermes 2.5 dataset by replacing Airoboros 2.2 with Airoboros 3.2.999https://huggingface.co/datasets/jondurbin/airoboros-3.2 Additionally, we incorporate the Aya dataset Singh et al.
+In this context, we select the OpenHermes 2.5 data collection Teknium (2023) as our base dataset, which is composed of multiple datasets covering a wide range of instructions and yielding excellent results when fine-tuned with Mistral-7B-v0.1.
+We make modifications to the original OpenHermes 2.5 dataset by replacing Airoboros 2.2 with Airoboros 3.2.888https://huggingface.co/datasets/jondurbin/airoboros-3.2 Additionally, we incorporate the Aya dataset Singh et al.
 (2024) to enhance the multilingual capabilities of our base model.
@@
 To bolster the model’s proficiency in Chinese, we include the COIG-CQIA Bai et al.
-(2024), ruozhiba-gpt4101010https://huggingface.co/datasets/hfl/ruozhiba_gpt4, and in-house Chinese multidisciplinary instruction data as supplementary datasets.
+(2024), ruozhiba-gpt4, [^9] https://huggingface.co/datasets/hfl/ruozhiba_gpt4 and in-house Chinese multidisciplinary instruction data as supplementary datasets.
 To enhance math and coding abilities, we use the dart-math-hard Tong et al.
@@
 (2023) datasets.
+The involved languages in the supervised fine-tuning stage can be found in Section C.2.
 Models m-ARC m-Hellaswag m-MMLU XWinograd XCOPA XStoryCloze
@@
 Llama-2-7B 35.5 48.6 35.4 78.0 58.9 55.6
-Mistral-7B-v0.1 40.7 54.5 46.7 80.5 55.8 60.2
+Mistral-7B-v0.1 40.7 54.5 46.7 80.5 55.8 57.2
 BLOOM-7B1 31.8 43.4 27.1 70.0 56.9 58.2
@@
 (2023) to train the DPO model.
-We leave the settings of post-training in Appendix A.
+We detail the settings of post-training in Section C.2.
 5 Experiments
@@
 LLaMAX2-Alpaca-7B 38.7 52.5 35.4 77.4 56.6 62.0 29.1 0.3
-FuxiTranyu-8B-SFT 31.8 51.5 26.8 75.7 61.3 56.6 25.9 8.9
-FuxiTranyu-8B-DPO 32.8 52.2 27.3 74.1 62.1 56.9 26.4 7.3
+FuxiTranyu-8B-SFT 32.8 49.2 26.9 74.7 61.2 57.4 28.3 9.2
+FuxiTranyu-8B-DPO 34.2 47.9 27.4 69.1 61.8 57.6 26.8 7.1
 Table 3: Average performance of FuxiTranyu-8B instruct and chat models compared to BLOOMZ-7B1, Llama-2-Chat-7B, and Mistral-7B-Instruct-v0.1 on mutlilingual discriminative and generative tasks.
-We conducted extensive experiments to evaluate the capabilities of FuxiTranyu under the multilingual setting, specifically from the base model to the instruction-tuned model.
-We selected several models as benchmarks to compare our models with both English-centric and multilingual models.
-For English-centric models, we compared FuxiTranyu against Llama-2 (Llama-2-7B, Llama-2-chat-7B) (Touvron et al., 2023b) and Mistral (Mistral-7B-v0.1, Mistral-7B-instruct-v0.1) (Jiang et al., 2023a).
+We conducted extensive experiments to evaluate the capabilities of FuxiTranyu (both the base model and instruction-tuned models) under the multilingual setting.
+We compared FuxiTranyu against strong baselines, including both English-centric and multilingual models.
+For English-centric models, we used Llama-2 (Llama-2-7B, Llama-2-chat-7B) (Touvron et al., 2023b) and Mistral (Mistral-7B-v0.1, Mistral-7B-instruct-v0.1) (Jiang et al., 2023a) as baseline.
 For multilingual models, we compared FuxiTranyu with BLOOM (BLOOM-7B1, BLOOMZ-7B1) (Scao et al., 2022a; Muennighoff et al., 2022), PolyLM (PolyLM-13B, PolyLM-MultiAlpaca-13B) (Wei et al., 2023), and LLaMAX2 (LLaMAX2-7B, LLaMAX2-7B-Alpaca) (Lu et al., 2024).
@@
 We evaluated the performance towards generative tasks, especially in translation and summarization tasks.
-For translation task, we employed WMT14 in en-fr translation direction (Bojar et al., 2014), WMT16 in en-de and en-ro translation directions (Bojar et al., 2016) and IWSLT 2017 (Cettolo et al., 2017) in en-ar translation direction for measuring the translation performance in our models and benchmark models.
-For summarization task, we used XL-Sum (Hasan et al., 2021) dataset.
+For the translation task, we employed WMT14 in en-fr translation direction (Bojar et al., 2014), WMT16 in en-de and en-ro translation directions (Bojar et al., 2016) and IWSLT 2017 (Cettolo et al., 2017) in en-ar translation direction for measuring the translation performance in our models and benchmark models.
+For the summarization task, we used XL-Sum (Hasan et al., 2021) dataset.
 We selected 15 languages for the evaluation (ar, en, es, fr, gu, hi, id, mr, pt, ru, sr, ta, uk, vi, zh).
 5.1 Base Model Evaluation
-First, we report experiment results of our base models vs.
+First, we report the experiment results of our base models vs.
 baseline models.
@@
 Evaluation results are shown in Table 2.
-Our model achieves the best performance on the XCOPA task.
+Our model achieves the best performance on the XCOPA and XStoryCloze tasks.
 For other tasks, our model is significantly better than multilingual models like BLOOM-7B and PolyLM-13B.
-When compared to LLaMAX-7B, the evaluation results of our model are almost comparable, with no significant difference from the evaluation results of LLaMAX-7B.
-But compared with english-centric models, our model still worse than Llama-2-7B and Mistral-7B-v0.1 due to the limited training data used for English.
+When compared to LLaMAX2-7B, the evaluation results of our model are almost comparable, with no significant difference from the evaluation results of LLaMAX2-7B.
+But compared with English-centric models, our model is still worse than Llama-2-7B and Mistral-7B-v0.1 due to the limited training data used for English.
 5.2 Instruction-Tuned Model Evaluation
-We further compared our instruction-tuned model with other instruction-tuned models.
+We further compared our instruction-tuned models with other instruction-tuned models.
 We evaluated these models on both discriminative and generative tasks.
@@
 On discriminative tasks, our models achieve the best result on XCOPA.
-For m-Hellaswag, XWinograd, and XStoryCloze, our models outperforms the English-centric models, but slightly underperforms the multilingual models compared with LLaMAX2-7B.
-Our models still underperforms in m-ARC and m-MMLU tasks due to the limited training data used.
-In generative tasks, our model excells on the summarization task, outperforming all baseline models.
-For the translation task, our model outperforms the English-centric models, but slightly underperforms the multilingual model like LLaMAX2-Alpaca-7B.
-More details of our evaluations are discussed in Appendix B, where we report the results for each language tested.
+For m-Hellaswag, XWinograd, and XStoryCloze, our models outperform the English-centric models but slightly underperform the multilingual models compared with LLaMAX2-7B.
+Our models still underperform in m-ARC and m-MMLU tasks due to the limited training data used.
+In generative tasks, our models excel on the summarization task, outperforming all baseline models.
+For the translation task, our models outperform the English-centric models but slightly underperform the multilingual model like LLaMAX2-Alpaca-7B.
+More details of our evaluations are discussed in Appendix D, where we report the results for each language tested.
 6 Analysis and Interpretability
@@
 (2024) perspectives.
-Specifically, our neuron analysis explores the importance of different neurons to multilingual abilities of the model, while the representation analysis examines the characteristics of multilingual representations learned by the model.
-Here, we first introduce the details and results of our neuron analysis, while the representation analysis is discussed in Section 6.2.
+Specifically, our neuron analysis explores the importance of different neurons to the multilingual abilities of the model, while the representation analysis examines the characteristics of multilingual representations learned by the model.
+Here, we first introduce the details and results of our neuron analysis, while the representation analysis is discussed in Section E.1.
 6.1 Neuron Analysis
@@
 The set $\bm{\alpha}$ can represent a component or a layer of the model, with the neuron indices in $\bm{\alpha}$ generally being continuous.
-6.1.1 Analysis Setup
+6.2 Neuron Analysis Setup
 We chose the Flores-200 dataset (Costa-jussà et al., 2022) to evaluate the importance of neurons.
 By selecting the languages ar, bn, es, fr, id, pt, ta, vi, zh, en, de, hu, it, ru, and sk, we analyzed the significance of different model components and layers in response to various linguistic inputs.
-6.1.2 Results
-We analyzed the varying importance of different layers across diverse language inputs, as shown in Figure 4 (Appendix C).
-Our findings indicate that, universally, shallow layers exhibit low significance while deep layers demonstrate great importance.
+6.3 Neuron Analysis Results
+We analyzed the varying importance of different layers across diverse language inputs, as shown in Figure 4 (Section E.2).
+Our findings indicate that universally, shallow layers exhibit low significance while deep layers demonstrate great importance.
 Notably, languages such as bn and ta exhibit a notably diminished importance in deep layers compared to others, aligning with our evaluation results where these languages perform poorly.
 This discrepancy may stem from their relatively limited representation learning in the pre-training data.
-We then analyzed the significance of various components across different language inputs, depicted in Figure 5 (Appendix C), with 8 components per layer.
+We then analyzed the significance of various components across different language inputs, depicted in Figure 5 (Section E.2), with 8 components per layer.
 Our findings mirror previous conclusions: components in shallow layers exhibit low importance, whereas those in deep layers show high significance.
 Moreover, a more detailed observation reveals that MLP components hold greater importance in shallow layers, whereas attention components are more critical in deep layers.
-Figure 3: Similarity distribution of multilingual representations in the intermediate layers of BLOOM-7B1 and FuxiTranyu-8B, with languages sorted based on their percentages in the pre-training data.
-6.2 Representation Analysis
-Language models encode textual symbols into high-dimensional representations with rich semantic information.
-For a multilingual language model, due to parameter sharing mechanisms, it encodes textual symbols from different languages into a unified representation space.
-Furthermore, through multilingual joint training, the model learns multilingual representations, which encode the intrinsic characteristics of languages and the relationships between different languages.
-Here, we explore the multilingual characteristics of the model from the perspective of the multilingual representations it learns.
-Specifically, we calculate the similarity of representations across different languages.
-To quantitatively evaluate the similarity between different language representations, we choose cosine similarity for its simplicity and effectiveness.
-To mitigate the impact of semantic differences on our analysis, we collect multilingual text data from open-source parallel corpora.
-For a language $l$, we input its corresponding text data into the model and collect text representations from the last token of each respective text.
-We then compute the average of these text representations to obtain the language representation $\bm{v}_{l}$ for language $l$.
-Finally, we calculate the similarity between two language representations as $\text{sim}(l_{1},l_{2})=\frac{\bm{v}_{1}^{\top}\bm{v}_{2}}{\|\bm{v}_{1}\|\|\bm{v}_{2}\|}$.
-It’s important to note that we extract language representations and compute similarity across each layer of the model.
-6.2.1 Analysis Setup
-We selected the Flores-200 dataset (Costa-jussà et al., 2022) as our parallel data source, which includes 2009 sentences for each language.
-For the explored languages, we chose en, zh, de, fr, es, ru, it, pt, nl, pl, ja, vi, cs, tr, hu, el, sv, ro, uk, and hi, based on their highest language proportions in our pre-training data.
-For comparison, we also analyzed the BLOOM-7B1 model (Scao et al., 2022a).
-For this model, we considered en, zh, fr, es, ru, pt, nl, pl, ja, vi, cs, tr, hu, el, sv, ro, uk, hi, fi, and th.
-6.2.2 Results
-Figure 3 illustrates the similarities distribution of multilingual representations in the intermediate layers of two models, with languages ordered according to the amount of language resources.
-It is apparent that for the BLOOM-7B, lower multilingual representation similarities tend to occur between the top 10 languages with higher resource availability and the bottom 10 languages with lower resource availability.
-In contrast, our model learn more consistent multilingual representations for all the languages we explored.
-This indicates that our model possesses a higher degree of multilingual balance, which is also reflected in our multilingual evaluation results and pre-training corpus.
-Furthermore, we calculate the average similarity for each layer of the two models, as shown in Figure 6 (Appendix C).
-For our model, it can be observed that there is a significant increase in similarity from the embedding layer to layer 0, reaching a very high level.
-As the depth of the model increases, the similarity continues to rise, indicating that the model learns richer multilingual alignment information in these layers.
-Subsequently, there is a sharp decrease in similarity from layer 28 to layer 29, suggesting that language-specific multilingual representations in the final layer are learned to predict the diverse multilingual vocabulary.
-For BLOOM-7B1, the trend of similarity changes across layers is similar, initially increasing and then decreasing, but the changes are more gradual in magnitude.
 7 Conclusion
-In this paper, we have presented the FuxiTranyu models to address the need for open-source multilingual LLMs.
-Along with the base model, FuxiTranyu-8B, we also present the fine-tuned models on multilingual supervised fine-tuning dataset and preference dataset, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO.
-Evaluations on multilingual benchmarks show FuxiTranyu models outperform previous multilingual and monolingual LLMs.
+In this paper, we have presented FuxiTranyu to address the need for open-source multilingual LLMs.
+Along with the base model, FuxiTranyu-8B, we also present instruction-tuned models fine-tuned on multilingual supervised fine-tuning and preference data, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO.
+Evaluations on multilingual benchmarks show FuxiTranyu outperforms previous multilingual and monolingual LLMs.
 Furthermore, interpretability analyses underscore the efficacy of the multilingual capabilities embedded in FuxiTranyu.
@@
 2023YFE0116400).
-The computing resources used in this project are supported by the Scientific Computing Center of CIC, Tianjin University.
-Appendix A Post-Training Details
-During the instruction tuning phase, we executed the fine-tuning process on 5 A100 80GB GPUs, leveraging the TRL framework for instruction fine-tuning and DPO training.
+The computing resources used in this project were supported by the Scientific Computing Center of the College of Intelligence and Computing, Tianjin University.
+We would like to thank the anonymous reviewers for their insightful comments.
+Appendix A Supported Languages in FuxiTranyu
+Our language selection strategy primarily stems from two distinct perspectives: the availability of pre-training data and geographical considerations.
+We initially approach language selection from the perspective of available pre-training data.
+Given that the majority of our pre-training data is sourced from web documents, e.g., CulturaX, we determine the languages for pre-training FuxiTranyu based on the statistical information derived from CulturaX.
+We select the top 21 languages based on the number of available tokens in descending order.
+Subsequently, we manually incorporate Asian languages, encompassing those from Southeast Asia, West Asia, and Central Asia, resulting in a total of 43 languages.
+The complete list can be found in Table 4.
+In terms of programming languages, we initially consider all 13 languages included in BLOOM Scao et al.
+(2022a), such as Java, JavaScript, and Python.
+Additionally, we include three programming languages (SQL, Assembly, and Visual Basic) due to their high popularity, as indicated by the TIOBE index.
+[^12] https://www.tiobe.com/tiobe-index/ The complete list of programming languages is provided in Table 5.
+Figure 2: Fertility test results of the tokenizers for FuxiTranyu, Llama-2, and BLOOM.
+Appendix B Tokenization
+We implement the Byte-level Byte-Pair Encoding (BBPE) algorithm using the Hugging Face tokenizer library.
+Our tokenizer is initiated from GPT-2’s tokenizer, incorporating both pre-tokenization and post-tokenization processes.
+Notably, we opt not to split numbers into digits.
+In line with the approach outlined in BLOOM Scao et al.
+(2022a), we expand the vocabulary size to 250,680 to accommodate multilingual scenarios, thereby mitigating the risk of over-segmentation in low-resource languages.
+For training the tokenizer, we randomly sample 1 million documents for each language from our collected data.
+It’s worth noting that for languages with a total document count being less than 1 million, we utilize all available documents in the training data for the tokenizer.
+Following the approach used in BLOOM, we also evaluate the performance of our tokenizer using the fertility metric.
+To assess its efficacy, we conduct a comparative analysis with the Llama-2 and BLOOM tokenizers.
+This evaluation involves computing fertility on the same set of documents across different languages.
+Results are presented in Figure 2, which indicate that the FuxiTranyu tokenizer is more efficient than the others in most languages.
+Based on our evaluations and interpretability analysis, we believe that the fertility of the tokenizer positively correlates with the model’s performance in specific languages.
+In the fertility test, we observe that Bengali (bn), Hindi (hi), and Tamil (ta) exhibit high fertility, indicating lower tokenization efficiency in these languages compared to others.
+Consequently, the performance and importance of neurons of these languages in our base model are also suboptimal.
+Further details are discussed in Section 6.3.
+Appendix C Training Details
+Pre-Training SFT DPO
+# Params 8B Learning Rate 2e-5 Learning Rate 5e-7
+Hidden Size 4,096 Warmup Ratio 10% Warmup Ratio 10%
+Intermediate Size 16,384 Batch Size 512 Batch Size 512
+Heads 32 Epochs 2 Epochs 1
+Layers 30 NEFTune $\checkmark$ beta 0.01
+FlashAttn V2 $\checkmark$ FlashAttn V2 $\checkmark$ FlashAttn V2 $\checkmark$
+Training Tokens 606B # Instances 1M # Instances 61.1k
+Position Embed 4,096
+Vocab Size 250,752
+Learning Rate 3e-4 $\rightarrow$ 1e-4
+Batch Size 2M $\rightarrow$ 4M
+Context Length 4,096
+Table 6: Model size and hyper-parameters. We append 72 dummy tokens to the vocabulary to make the embedding size be divisible by 128.
+C.1 Pre-training Details
+The training procedure for the FuxiTranyu model adheres to the standard autoregressive language model framework, utilizing the next-token prediction loss as detailed in (Brown et al., 2020).
+To enhance pre-training efficiency, we employ a document packing method similar to that described in (Raffel et al., 2020).
+This involves randomly shuffling documents, merging them, and then truncating them into multilingual chunks that adhere to a maximum context length of 4096 tokens during the pre-training phase.
+To mitigate memory consumption and further improve training efficiency, we leverage ZeRO-2 Rajbhandari et al.
+(2020) and Flash-Attention V2 Dao (2024) technologies.
+For optimization, the standard AdamW optimizer Loshchilov and Hutter (2017) is utilized with hyper-parameters set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$.
+We employ the cosine learning rate scheduler, starting with a maximum learning rate of 3e-4 and decaying to a minimum of 10% of the maximum rate.
+Notably, after encountering divergence issues when training approximately 241 billion tokens, we reduced the maximum learning rate to 1e-4 to match with the learning rate used in BLOOM, given the multilingual context of both models.
+Our FuxiTranyu-8B model is trained using the Megatron-LM Shoeybi et al.
+(2019) framework on a setup of 32 A800 GPUs, processing a total of 606 billion tokens.
+The training utilizes FP16 mixed precision to ensure stability.
+Detailed training parameters and configurations are provided in Table 6.
+C.2 Post-Training Details
+The instruction datasets collected do not cover all languages used during pre-training.
+For the current version of FuxiTranyu, we provide support for the following languages: Arabic, Bengali, Burmese, Chinese, Dutch, English, Filipino, Finnish, French, German, Greek, Hindi, Hungarian, Indonesian, Italian, Japanese, Korean, Kyrgyz, Malay, Persian, Polish, Portuguese, Russian, Spanish, Swedish, Tamil, Thai, Turkish, Ukrainian, Urdu, and Vietnamese.
+During the instruction tuning phase, we executed the fine-tuning process on 8 A800 80GB GPUs, leveraging the TRL framework for instruction fine-tuning and DPO training.
 Throughout both stages, we employed the ChatML format [^13] https://github.com/openai/openai-python/blob/release-v0.28.0/chatml.md for the chat template, and designated <PAD> as the pad token.
@@
 In the SFT stage, we configured the maximum learning rate to 2e-5, with a warmup phase spanning 10% of the total steps.
-The global batch size was set to 320, and the model was trained for 2 epochs.
+The global batch size was set to 512, and the model was trained for 2 epochs.
 To optimize memory usage, we enabled Flash-Attention V2 Dao (2024), ZeRO stage 2 Rajbhandari et al.
@@
 The maximum learning rate was set to 5e-7, with a warmup phase covering 10% of the total training steps.
-Similar to the SFT stage, the global batch size was maintained at 320, and we activated Flash-Attention V2 and gradient checkpointing to optimize memory usage.
+Similar to the SFT stage, the global batch size was maintained at 512, and we activated Flash-Attention V2 and gradient checkpointing to optimize memory usage.
 To accommodate the policy and reference model within memory constraints, we utilized ZeRO stage 3 for the policy model and omitted ZeRO for the reference model.
-Appendix B Detailed Evaluation Results
+Appendix D Detailed Evaluation Results
 We provide detailed evaluation results for each language in this section.
 First, we present the results for all 15 tested languages on the multilingual ARC in Table 7, comparing base models and instruction-tuned models.
-The results show that our models perform better in 1 of the 15 tested languages for the ARC task.
-We speculate that our models still underforms on this task due to the relatively small amount of training data used.
+In base models, the results show that our models perform better in 1 of the 15 tested languages for the ARC task.
+In instruction-tuned models, our models outperforms in ar and vi languages.
+We speculate that our models still underperformed on this task due to the relatively small amount of training data used.
 Models ar bn de en es fr hu id
@@
 […]
-Instuction-tuned Model
+Instruction-tuned Model
 Llama-2-Chat-7B 26.2 23.9 39.8 53.6 43.0 42.5 32.4 35.4
 […]
-[…]
+BLOOMZ-7B1 31.2 26.2 25.4 42.7 37.2 37.6 22.8 35.9
 PolyLM-MultiAlpaca-13B 27.4 18.4 30.5 38.2 32.9 32.8 18.6 30.2
 […]
+FuxiTranyu-8B-SFT 31.1 26.3 33.9 38.9 35.4 36.3 31.5 35.1
 […]
-FuxiTranyu-8B-DPO 32.4 26.9 33.8 36.3 35.3 35.5 34.0 33.7
 Models it pt ru sk ta vi zh
@@
 […]
-Instuction-tuned Model
+Instruction-tuned Model
 […]
@@
 […]
-FuxiTranyu-8B-SFT 33.7 33.3 31.1 28.2 23.4 31.9 34.6
-FuxiTranyu-8B-DPO 34.6 34.2 32.5 29.3 24.6 32.5 36.9
+FuxiTranyu-8B-SFT 33.6 35.2 32.2 29.0 23.5 32.5 36.8
+FuxiTranyu-8B-DPO 36.8 37.1 33.8 29.1 25.1 33.7 37.4
 Table 7: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B, PolyLM-13B, and LLaMAX2-7B models on multilingual ARC (25-shot).
@@
 Despite our FuxiTranyu-8B model being trained on only about 600B tokens, it achieves remarkable performance.
-The SFT and RL-trained models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also deliver promising results across all languages, even competing with powerful monolingual LLMs like Llama-2-7B and Mistral-7B-v0.1, with English language as exception.
+Comparing base models, our models outperforms other models in ar, bn, hu, and vi languages.
+Our model still lag behind compared with Mistral-7B-v0.1, but outperform other baseline models, except sk language.
+The SFT and RL-trained models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also deliver promising results across all languages, even competing with powerful monolingual LLMs like Llama-2-7B and Mistral-7B-v0.1, with English and Spanish as an exception.
 Models ar bn de en es fr hu id
@@
 PolyLM-13B 39.6 28.4 49.5 71.3 55.8 54.8 29.3 50.1
-LLaMAX2-7B 43.3 32.3 53.8 75.4 59.0 58.1 44.1 51.0
 […]
-Instuction-tuned Model
-Llama-2-Chat-7B 31.4 28.3 50.7 78.6 58.1 57.0 39.0 44.5
-Mistral-7B-Instruct-v0.1 31.2 28.7 52.2 70.1 58.1 57.6 39.8 38.1
+[…]
+Instruction-tuned Model
+[…]
+[…]
 BLOOMZ-7B1 39.5 31.5 33.1 46.6 48.7 45.7 29.8 42.0
@@
 […]
-FuxiTranyu-8B-DPO 48.1 33.6 57.7 57.8 62.5 62.5 49.3 54.5
+[…]
 Models it pt ru sk ta vi zh
@@
 […]
-Instuction-tuned Model
+Instruction-tuned Model
 Llama-2-Chat-7B 53.7 54.0 47.6 36.4 28.8 41.2 45.1
@@
 BLOOMZ-7B1 40.3 37.3 33.1 29.6 29.5 40.6 42.6
-[…]
+PolyLM-MultiAlpaca-13B 40.8 42.4 40.0 27.1 25.2 38.2 53.5
 […]
@@
 We report results on multilingual MMLU in Table 9.
-Our models still underperforms baseline models for all languages.
+Our models still underperform baseline models for all languages.
 It is in line with the number of training tokens utilized in the pre-training process.
@@
 FuxiTranyu-8B 26.3 25.5 27.6 27.1 27.1 27.5 26.4 26.2
-Instuction-tuned Model
+Instruction-tuned Model
 […]
@@
 LLaMAX2-7B-Alpaca 30.0 30.4 36.4 43.0 37.2 36.9 47.6 35.5
-FuxiTranyu-8B-SFT 26.0 27.1 26.6 27.0 26.4 27.8 27.3 26.3
-FuxiTranyu-8B-DPO 27.0 27.3 27.2 27.0 27.4 27.8 27.6 26.4
+FuxiTranyu-8B-SFT 26.2 26.8 27.5 28.2 28.1 27.6 26.0 25.9
+FuxiTranyu-8B-DPO 27.3 27.8 27.7 28.2 27.9 27.3 26.8 26.6
 Models it pt ru sk ta vi zh
@@
 FuxiTranyu-8B 27.1 26.8 27.7 26.0 26.3 26.3 26.0
-Instuction-tuned Model
+Instruction-tuned Model
 […]
@@
 […]
-FuxiTranyu-8B-SFT 27.1 27.0 26.8 27.2 26.4 25.9 27.0
-FuxiTranyu-8B-DPO 27.5 27.7 28.0 27.6 26.9 26.2 27.7
+FuxiTranyu-8B-SFT 26.2 25.9 27.9 26.6 27.0 26.4 26.8
+FuxiTranyu-8B-DPO 28.6 27.1 28.2 26.7 26.8 26.7 28.0
 Table 9: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B, PolyLM-13B, and LLaMAX2-7B models on multilingual MMLU (5-shot).
 Results on XWinograd are depicted in Table 10.
-Our FuxiTranyu SFT and DPO models achieve better results in Portuguese and Chinese.
-Although our models underperforms in English, French, Russian, and Japanese compared to Llama-2-7B, they outperforms previous multilingual LLMs like BLOOM-7B1 and PolyLM-13B across all languages.
+In base models, although our models still underperformed compared to Mistral-7B-v0.1, they outperforms previous multilingual LLMs like BLOOM-7B1 and PolyLM-13B across all languages.
+Notably, our FuxiTranyu SFT and DPO models achieve better results in Chinese.
 Models fr pt zh en ru jp
@@
 […]
-Instuction-tuned Model
+Instruction-tuned Model
 […]
@@
 […]
-[…]
+FuxiTranyu-8B-DPO 77.1 67.3 66.7 73.9 62.9 66.5
 Table 10: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XWinograd (5-shot).
@@
 For XCOPA, our base models achieve better results in sw, ta, tr, and vi.
-When compared to instruction-tuned models, our models achieve better results in more languages, specifically in it, id, ta, th, tr, vi, and zh.
+When compared to instruction-tuned models, our models achieve better results in more languages, specifically in it, ta, th, tr, and vi.
 On the XStoryCloze task, our base models achieve better results in three languages: ar, my, and ru.
-However, for instruction-tuned models, our models outperforms other baseline models only in my.
+However, for instruction-tuned models, our models outperform other baseline models only in my.
 Models et ht it id qu sw ta th tr vi zh
@@
 […]
-Instuction-tuned Model
+Instruction-tuned Model
 Llama-2-Chat-7B 47.8 51.4 67.0 62.4 50.8 52.2 50.6 54.8 55.6 61.6 61.2
-Mistral-7B-Instruct-v0.1 48.2 51.2 65.4 54.0 49.2 54.6 55.2 53.2 52.2 53.2 63.4
+[…]
 […]
@@
 […]
-Instuction-tuned Model
+Instruction-tuned Model
 […]
@@
 We present our evaluation results for generative tasks in Table 13 and Table 14.
-On the XL-Sum task, our models significantly outperform all baseline models across all evaluated languages, demonstrating the potential of our models for summarization task, particularly in a multilingual context.
-For the translation tasks in WMT14, WMT16, and IWSLT2017, our models excell in the en-ro, en-de, and en-fr translation directions.
+On the XL-Sum task, our models significantly outperform all baseline models across all evaluated languages, demonstrating the potential of our models on summarization task, particularly in a multilingual context.
+For the translation tasks in WMT14, WMT16, and IWSLT2017, our models excel in the en-ro and en-de translation directions.
 However, they still lag behind other baseline models in the ro-en, de-en, fr-en, ar-en, and en-ar translation directions.
 This indicates that our models perform significantly better for out-of-English translation directions.
-Although our models underperform in the en-ar direction compared to LLaMAX-2-Alpaca, they still achieve notably better results than other models.
+Although our models underperformed in the en-fr and en-ar directions compared to LLaMAX2-Alpaca, they still achieve notably better results than other models.
 Models ar en es fr gu hi id mr pt ru sr ta uk vi zh
 […]
-Mistral-7B-Instruct-v0.1 0.1 11.0 3.0 3.4 0.3 0.2 3.1 0.6 3.2 0.4 2.1 0.2 0.3 4.6 0.6
+[…]
 […]
@@
 […]
-Table 13: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XL-Sum (0-shot).
+Table 13: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, and LLaMAX2-7B models on XL-Sum (0-shot).
 Models WMT16 (EN-RO) WMT16 (RO-EN) WMT16 (EN-DE) WMT16 (DE-EN)
 BLEU CHRF BLEU CHRF BLEU CHRF BLEU CHRF
-[…]
+Llama-2-Chat-7B 17.18 44.20 31.43 58.00 20.01 48.31 35.41 60.78
 Mistral-7B-Instruct-v0.1 13.66 41.47 24.58 53.04 19.41 49.25 30.19 58.27
 BLOOMZ-7B1 1.88 20.09 11.35 36.22 3.76 23.27 22.30 46.69
-LLaMAX2-7B-Alpaca 24.52 51.94 36.02 60.85 26.31 53.95 37.05 61.90
+[…]
 […]
@@
 Mistral-7B-Instruct-v0.1 24.24 52.08 31.40 59.50 9.13 32.64 0.31 13.31
+BLOOMZ-7B1 17.73 41.02 31.07 56.03 25.25 47.64 4.58 25.05
 […]
-LLaMAX2-7B-Alpaca 32.86 59.53 36.00 61.64 29.76 52.68 10.47 40.27
-FuxiTranyu-8B-SFT 34.06 60.74 28.83 57.86 21.42 42.91 8.19 35.67
 […]
-Table 14: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on WMT14, WMT16, and IWSLT2017 (0-shot).
-Appendix C Detailed Analysis Results
+FuxiTranyu-8B-DPO 31.98 59.64 32.27 60.19 27.05 51.5 6.5 29.41
+Table 14: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, and LLaMAX2-7B models on WMT14, WMT16, and IWSLT2017 (0-shot).
+Appendix E Additional Analysis and Interpretability
+Figure 3: Similarity distribution of multilingual representations in the intermediate layers of BLOOM-7B1 and FuxiTranyu-8B, with languages sorted based on their percentages in the pre-training data.
+E.1 Representation Analysis
+Language models encode textual symbols into high-dimensional representations with rich semantic information.
+For a multilingual language model, due to parameter sharing mechanisms, it encodes textual symbols from different languages into a unified representation space.
+Furthermore, through multilingual joint training, the model learns multilingual representations, which encode the intrinsic characteristics of languages and the relationships between different languages.
+Here, we explore the multilingual characteristics of the model from the perspective of the multilingual representations it learns.
+Specifically, we calculate the similarity of representations across different languages.
+To quantitatively evaluate the similarity between different language representations, we choose cosine similarity for its simplicity and effectiveness.
+To mitigate the impact of semantic differences on our analysis, we collect multilingual text data from open-source parallel corpora.
+For a language $l$, we input its corresponding text data into the model and collect text representations from the last token of each respective text.
+We then compute the average of these text representations to obtain the language representation $\bm{v}_{l}$ for language $l$.
+Finally, we calculate the similarity between two language representations as $\text{sim}(l_{1},l_{2})=\frac{\bm{v}_{1}^{\top}\bm{v}_{2}}{\|\bm{v}_{1}\|\|\bm{v}_{2}\|}$.
+It’s important to note that we extract language representations and compute similarity across each layer of the model.
+E.1.1 Analysis Setup
+We selected the Flores-200 dataset (Costa-jussà et al., 2022) as our parallel data source, which includes 2009 sentences for each language.
+For the explored languages, we chose en, zh, de, fr, es, ru, it, pt, nl, pl, ja, vi, cs, tr, hu, el, sv, ro, uk, and hi, based on their highest language proportions in our pre-training data.
+For comparison, we also analyzed the BLOOM-7B1 model (Scao et al., 2022a).
+For this model, we considered en, zh, fr, es, ru, pt, nl, pl, ja, vi, cs, tr, hu, el, sv, ro, uk, hi, fi, and th.
+E.1.2 Results
+Figure 3 illustrates the similarities distribution of multilingual representations in the intermediate layers of two models, with languages ordered according to the amount of language resources.
+It is apparent that for the BLOOM-7B, lower multilingual representation similarities tend to occur between the top 10 languages with higher resource availability and the bottom 10 languages with lower resource availability.
+In contrast, our model learn more consistent multilingual representations for all the languages we explored.
+This indicates that our model possesses a higher degree of multilingual balance, which is also reflected in our multilingual evaluation results and pre-training corpus.
+Furthermore, we calculate the average similarity for each layer of the two models, as shown in Figure 6 (Section E.2).
+For our model, it can be observed that there is a significant increase in similarity from the embedding layer to layer 0, reaching a very high level.
+As the depth of the model increases, the similarity continues to rise, indicating that the model learns richer multilingual alignment information in these layers.
+Subsequently, there is a sharp decrease in similarity from layer 28 to layer 29, suggesting that language-specific multilingual representations in the final layer are learned to predict the diverse multilingual vocabulary.
+For BLOOM-7B1, the trend of similarity changes across layers is similar, initially increasing and then decreasing, but the changes are more gradual in magnitude.
+E.2 Detailed Analysis Results
 We present the varying importance of different layers across diverse language inputs in Figure 4.